A Bibliography of Publications of John A. Nelder

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

11 April 2026 Version 1.294

Abstract

This bibliography records publications of John A. Nelder.

Title word cross-reference

2^{n} [Nel63b]. ^{2} [Baş26]. H [Boo07, Dry10, LNN07, Rob07].

-Likelihood [Rob07, Boo07, Dry10, LNN07].

12th [Jef72]. 14th [BvdH00]. 16 [Nel98b]. 18th [BS86]. 19 [Nel01b]. 1971 [Jef72]. 1976 [Ano75, BS^{+}97]. 1978 [CH78, Fos79]. 1979 [DLPT80]. 1983 [Did84]. 1984 [Bil85]. 1986 [BS86]. 1987 [FU88]. 1991 [GH92].

2.5 [Fos79]. 20 [Nel69b, Nel94c]. 2000 [BvdH00]. 2010 [Sen19]. 2nd [Che90].

3 [Col79]. 3289 [Nel01b]. 3rd [CH78].

8 [Nel68a].

9 [Nel68b]. 94g [Nel94c]. 96 [Nel76b]. 9th [Ano75].

A. [Les97, LR97, Nel97c, Wir97]. Aachen [Hau86]. ab-initio} [PP23]. Acceleration [YCH24]. Accessing [LIN88]. accurate [CS03]. acoustics [KNCS21]. Adams [Ano07]. adaptive [GH12, Meh20, WYWW15]. Additive [Nel68b, Nel86a]. aerobic [GGN64]. Afifi [Nel79b]. aggregates [GGN64]. agricultural [Nel67]. alas [Nel97b]. Algorithm [CGVP15, Gal24, Kel99, LPW12, Nel76b, TFB21, DW87, YCH24, Arn93, AT18, Baş26, BPT06, FBP19, GH12, KBČ20, KNCS21, LTJ95, LL12, MM26, Meh20, MJ10, NZE^{+}15, NT02, Nel80c, PCB02, PB26, RWJ13, SGK16, SS04, SEABZ24, Tom95, TA13, WYWW15, WSKF13, ZZP^{+}22, Nel68b, Nel68a, Nel69b, Nel76b]. Algorithms [Nel68b, Nel68a, Nel69b, NCC^{+}70, Nel76b, Pay04, CS03, CS05, Mas06, Mas07, Mas09a, Mas09b, SGK16]. alive [Nel97b]. Alternative [Nel62a, Nel85c, Nel94a, LN06b]. Altman [Nel01b]. Amongst [Rip04]. analysing [NL01]. Analysis [Boo07, DLPT80, DCZ^{+}19, LN82, LN02, Nel75b, Nel79d, Nel79c, Nel79a, Nel79b, Nel82a, Nel87c, NGMS94, NL95, Nel00e, Rob07, ANB64, BC97, Did84, Dry10, HNX03, LN98, LN00a, Nel51, Nel62c, Nel65a, Nel65b, NL91, Nel94c, Nel95a, Nel98b, LN99b, Pay04, SS01, WON88, Nel72c, Nel78a, Nel72b]. Analytical [CCG12]. annealing [TA13]. Announcement [Nel75a]. ANOVA [CST04]. Anwendung [FU88]. Application [IBV^{+}22, Nel77a, OT09, BC97, KBČ20, LN00b, MM26, Nel67, Nel98b, NMD^{+}98, Nel56a]. approach [NABS60, PP23, Nel79b]. areas [Nel93b]. Argumented [YCH24]. arising [HNX03]. Arithmetic [Aki77]. Array [Nel69b]. arrays [Mas07]. Artificial [Nel87a, Phe87, Str88]. ascent [Arn93]. Ashworth [Ano11, MGP11, Sen19]. Aspects [Nel93a]. assessment [Nel72e]. assistant [Nel92a]. Atlanta [Bil85]. August [Ano75, Sen19]. Austria [Fos79]. Automatique [Did84]. Azen [Nel79b].

B [Nel79a, Nel87b, Nel88c]. B. [Nel82b]. back [Ait95, Gow95, Nel94d, Nel95b, Nel95c, RTW95, Sea95, Van95]. Baden [Fos79]. Baker [Alt80, Col79]. Balance [Sen04]. balanced [Nel68d]. band [DCZ^{+}19]. barrier [Pri21]. based [AT18, Baş26, DCZ^{+}19, Lee04, LL12, Meh20, NT02, Nel91, WYWW15, WON88, XD14, ZZP^{+}22]. basic [Van95]. basics [Nel94d, Nel95c, Ait95, Gow95, Nel95b, RTW95, Sea95]. basis [WYWW15]. bat [SEABZ24]. bat-Nelder [SEABZ24]. Bayesian [Nel98b, BC97]. Behavioural [Nel79c]. Besag [Ano11]. between [Arn93, LN00b]. Beyond [BJM14, Lee04]. bilevel [RWJ13]. biologists [Nel72d]. Biology [Nel75f, Nel71a]. Biometric [Ano75, Hau86]. biometrical [Nel52, Nel53]. Biometry [Nel63a]. Bliss [Nel71a]. Block [Cal71, Nel65a, Nel65b, Nel65a]. BMDP [Nel78c]. Book [Alt80, Ano07, Boo07, Bur91, Che90, Col79, Dal72, Dry10, Gar71, Gho08, Inn70, K.71, Lac90, Nel56a, Nel63a, Nel64a, Nel68c, Nel71a, Nel72b, Nel72c, Nel75b, Nel77a, Nel78a, Nel79d, Nel79c, Nel79a, Nel79b, Nel80b, Nel80a, Nel87b, Nel94b, Pla84, Rob07, Sch92, Wat86, Wil84, Nel82b]. Boston [Ano75]. Both [Sen04]. boundary [Mas09a]. bounded [WSKF13]. Breeding [Nel62c]. Brian [Nel90b]. British [Jef72]. Broemeling [Nel98b]. Building [Nel68e, Nel59].

C [Dal72, Gar71, Inn70, K.71, Nel64a]. calibration [WSKF13]. carrots [ANB64]. Categorical [Nel75b]. centroid [FBP19]. Century [JK97]. Certain [Nel72b, HNX03, Nel60]. Chambers [Nel78a]. Characteristics [CCG12]. Chatterjee [Nel79a]. Chebyshev [Meh20]. Chester [Nel71a]. Chronicle [BJM14]. circuits [NMS98]. cities [KNCS21]. class [Nel98d, RWJ13]. Classes [Rip04]. classical [Nel74b]. clinical [Lar97]. cliniques [Nel64a]. code [PP23]. coded [MM26]. Coefficient [Nel94b]. cold [NM56a]. Collins [BS86]. Colorado [BS86]. combination [Nel68d]. combinations [Lar97]. combining [CS05]. comme [Nel80d]. Comment [Car97, Nel78c, Nel86a, Nel88c, Wir97, Nel94c]. Comments [Ait95, Gow95, RTW95, Sea95, Nel95b, Van95]. Comparison [TFB21, Arn93]. comparisons [NL92]. Complexity [SS01]. Component [Nel71c]. components [CJM^{+}62, Nel52, Nel54a, Nel60]. COMPSTAT [BvdH00, BFS74, CH78]. Computation [Gol69, MN69, MN14, Dal72, Gar71, Inn70, K.71]. Computational [BvdH00, BFS74, CH78, Hau86, Pay04, Nel78a]. computationally [TOO23]. Computer [Nel71b, Nel79b, NL95, Nel92a, Bil85, BS86]. Computers [Nel75f, Nel67, Nel80b]. Computing [BT91, NC71, Nel71b, Nel75a, Nel84b, Nel69a, NP75, Nel76a, NP91, Nel93a, Nel80a]. Condition [Kel99]. Conditional [LN04]. Conference [Ano75, Fos79, GH92]. Conferences [Nel83b]. conjecture [Ham52, Uli58]. Conjugate [Nel94c]. Consensus [Paw04]. constants [IBV^{+}22]. Constrained [TFB21, AT18, Pri21]. constraints [Bre04, Lar97]. Construction [Nel68b, NCC^{+}70]. Contingency [Nel77b, Nel00e, Nel74b]. Continuous [DDM^{+}16, BC97, CS03, CS05, CJM^{+}62, Nel98b]. contrast [Sel15]. Contrasts [Nel63b]. Contribution [Nel97c]. control [Baş26]. controller [Baş26, LTJ95]. Controversies [Paw04]. Controversy [NLSS08]. Convergence [Gal22, Gal24, LRWW98, LRWW99, LPW12, McK98, McK99, SEABZ24, Tom95]. convergent [Cha12, PCB02]. Conversation [Sen03]. Cook [Nel98b]. Cooper [Nel90b]. coordination [LY12]. Correction [Ano61, ZZP^{+}22]. correlated [LN00a, NL01]. Correspondence [CNM00, Nel01a]. Corrigenda [Ait59]. Covariance [LN82]. Credibility [NV97]. Critical [Str88]. Criticism [Cox04]. crop [BN60, GFNF82, NABS60]. Cross [Nel79d]. Cross-Tabulated [Nel79d]. Crowder [Ano07]. crucial [Wes19]. cruise [Baş26]. Curve [Ano61, Nel61a]. curves [Nel92d].

D [Ano07, Nel64a, Nel98b]. dans [Nel80d]. Data [DLPT80, Did84, Nel68f, Nel69b, Nel72c, NP75, Nel78a, Nel79c, Nel85b, NRCC91, Nel94a, NGMS94, LN00a, LN00b, LN00c, LN02, Nel63d, NBCS66, Nel69a, Nel82d, Nel92e, Nel94c, Nel00d, NL01, Pay04, Nel75b, Nel79d]. d’Automatique [DLPT80]. Dave [Ano07]. DDE [SEABZ24]. December [Fos79, Hau86]. Decomposition [Nel85c, Mas07, WYWW15]. Decompositions [Gol69]. Deconstructing [NGL^{+}94]. Decrease [Kel99]. density [Nel63d, NM98]. derived [Nel98d]. Description [NCC^{+}70, Nel69a]. Design [NAM^{+}92, Nel56a, Nel72b, Baş26, Nel79e, NMS98, NMD^{+}98, NL04b]. designed [Bai04]. Designs [Cal71, Nel62b, Nel68d, Nel82b]. Desirable [Cal71, Nel80d]. Desmond [Car97, Les97, LR97, Nel97c, Wir97]. Detection [Kel99, LN03b]. Developments [Nel84b, BS^{+}97]. device [NMD^{+}98]. differentiable [DDM^{+}16]. differential [Mas06]. diffusion [GGN64]. Digital [Nel80b]. Dimensional [Nel77b, Arn93, FBP19]. dimensionality [HN06]. Dimensions [LRWW98, LPW12, LRWW99, Lóc13, Meh20]. d’Informatique [DLPT80]. direct [Cha12, KiANS22]. direction [XD14]. directional [LY12]. discretized [WSKF13]. Discussion [Les97, LR97, NAM^{+}92, Nel87c, Nel89, NGL^{+}94, Nel97c]. Dispersion [CCG12, NLB^{+}98, Nel92b, LN99b]. dispersions [LN01]. distinctions [Van95]. distributed [KTL19]. distribution [RWJ13]. distributions [Arn93, Nel94c]. diversification [SEABZ24]. Do [Nel84a]. dose [Lar97]. Double [LN06a, NL94, NL04a, LN00b]. double-exponential [LN00b]. Douglas [Nel01b]. Dr. [Nel90b]. drugs [GN64]. Dutch [NM56a].

each [Nel88a]. earthquake [TA13]. ECG [WYWW15]. Ecological [Jef72]. ecology [Jef72]. economic [SGK16]. Editor [BLN^{+}81, Nel98b, Nel01b, YYB^{+}76]. Editorial [Nel64b]. Editors [NS74, NKA92, NR97, Nel98c, NLSS08]. Edn [Che90]. Edward [Nel90b]. Effect [GGN64, HN06, GN64, LN01, LN05, LN06b]. Effects [Boo07, LNP06, LNP17, Nel68a, Rob07, ANB64, Dry10, NZE^{+}15, Gho08]. Efficient [Nel69b, SS04]. Eigenvalue [TFB21]. electromagnetic [NMD^{+}98]. electronic [NMS98]. elements [Mas09b]. Encyclopaedia [FKJ^{+}89]. End [NW86, Nel91, WN86, WON88]. enhanced [FZ06]. Enslein [Nel80b]. environment [NP91]. environments [GH92]. Enzyme [NRCC91]. Enzyme-Kinetic [NRCC91]. Epilogue [NC67]. Equation [Nel62a]. equations [Mas06, Mas09b, NL97]. Errata [NM65a]. error [Arn93]. Errors [NF84, BC97, Nel98b]. Essais [Nel64a]. Essence [Nel63a]. Estimating [CJM^{+}62, NGN92, TA13, Des97a, Les97, LR97, Nel97c, NL97, Wir97]. Estimation [Nel71c, GT89, Nel60, RWJ13]. estimator [LN99a]. estimators [LN03a]. European [BS^{+}97]. Evaluation [Nel74d, Fos79, Lar97, Nel79e]. Evolutionary [Nel66a]. examination [Nel62d, Nel04b]. Example [Nel79a]. executing [Nel56b]. expensive [TOO23]. experiment [Nel60]. Experimental [Nel79e, NMS98, Nel56a]. experimentation [Nel66a]. Experiments [Nel62b, Nel63b, NL95, Bai04, LN98, Nel56b, Nel63c, Nel65a, Nel65b, NL91, Nel95a, LN99b, Nel72b]. Expert [Hau86, Nel88b, Nel87a, Nel88a, Str88]. exponential [LN00b]. Extended [LN03a, NP87, NL97, KTL19, LN00b]. Extended-REML [LN03a]. extending [LIN88]. extension [Bre13, GT89].

F [LR97, Nel94c, Nel97c, Wir97]. F.R.G. [Hau86]. Factor [Nel66b, Nel77b, Nel00e]. Factorial [Nel82b, NL95]. Failure [NGMS94, Nel94c]. False [LN03b]. families [LN00b]. Fast [WYWW15]. Federer [Nel82b, Nel56a]. fertilizer [GFNF82]. Fiber [CCG12]. field [Nel56b, NBCS66]. find [Lar97]. finite [Mas09b]. Fitting [Ano61, LN06b, Nel61a, Nel00a, SMNS01]. Flamant [Nel64a]. flight [Baş26]. flight-based [Baş26]. flourishing [Nel97b]. flow [NZE^{+}15]. flux [NM98]. Folks [Nel72c]. Foreword [Nel04a]. Form [Nel62a]. Formation [Nel69b, Nel72a, NPT^{+}74, Nel75e]. Forms [Ait50, Ait59, Nel51]. Formulation [CCG12]. Fort [BS86]. Fortran [LIN88]. Fortschritte [FU88]. Fourier [WYWW15]. Fractional [Nel63b]. frames [NM56a]. Frank [NL95, Nel97a]. Franklin [Nel77a]. French [Nel80d]. frequencies [Nel52]. Front [NW86, Nel91, WN86, WON88]. Front-End [NW86]. FRS [ACHS04c, Ano07]. Function [Aki77, NP87, ON75, FBP19, NM65a, NM65b, NMD^{+}98]. Functional [NR97, Nel00a]. Functions [Nel66b, CS03, CS05, Des97a, Les97, LR97, Nel97c, Wir97]. future [Nel78b, Nel93a]. Fuzzy [LL12, LTJ95].

G [Nel79d, Nel01b]. gaps [DCZ^{+}19]. Gaussian [HN55]. Geissler [LN00b]. General [CST04, Nel79c, Sen04, AGL^{+}70, Nel65b, NG71]. general-purpose [NG71]. Generalised [Sen04, LN01]. Generalization [Ano61, MJ10, Nel61a, Nel74b]. Generalized [BN78, LN98, LNP06, LNP17, MN83, MN89, Nel62a, NW72, Nel75a, Nel82a, NB82a, Nel87a, NL91, NRCC91, Nel95a, LN96, LN02, LN06a, NGN92, NL94, NV97, Nel98d, NL04a, NL04b, Dry10, Nel82c, Nel86a, Pre84, Boo07, Bur91, Che90, Gho08, Lac90, Pla84, Rob07, Sch92, Wat86, Wil84]. generally [Nel68d]. Genetic [CS03, Mas09a, MM26, Mas06, Mas07, Mas09b, NZE^{+}15]. genetics [Nel52, Nel53, Nel60, Nel62c]. genotypic [Nel52]. GENSTAT [NP91, AGL^{+}70, LIN88, Nel74a, Nel75c, Nel80d]. Gentle [Nel80a]. Geometric [MJ10]. George [Nel94c]. Georgia [Bil85]. German [Hau86]. Germany [GH92]. Gilding [NT02]. Given [CGVP15]. glasshouses [NM56b]. GLIM [Alt80, Col79, BN78, Nel75a, Nel82a, NB82b, Nel85a, NW86, Nel87a, Nel91, WN86]. GLIMPSE [Nel91, WON88]. GLMs [LN03b, Lee04]. global [CS03, CS05, SEABZ24]. globally [Cha12]. Gnanadesikan [Nel72b]. Godambe [Nel89]. golden [NT02]. golden-section [NT02]. gradient [NMD^{+}98]. Graham [Nel79d]. Grange [Jef72]. Grange-over-Sands [Jef72]. Graphics [BT91]. Graphs [Nel76b]. Graybill [Nel77a]. great [Nel97b]. Grenoble [BS^{+}97]. Grid [BPT06]. Group [Hau86, Nel66b]. groups [Nel63c]. growth [ANB64, GN64, Nel61b, Nel62c]. Guide [Nel74d, NMT^{+}98].

H [Nel80b]. Hammersley [Dvo57]. Hand [Ano07]. Hardware [YCH24]. Harris [ZZP^{+}22]. Hastings [Nel62d, Nel04b]. hawks [ZZP^{+}22]. heated [NM56b]. heating [NZE^{+}15]. Hedayat [Nel82b]. Heidelberg [FU88]. held [BvdH00, CH78]. helpful [Nel00b]. Heredity [Nel98e]. HGLMs [LN00a, LNP11]. Hierarchical [LN96, LN01, Sen04, LN02, LN06a, NL04a]. high [FBP19, GH92, Meh20]. high-dimensional [FBP19]. high-level [GH92]. holders [NMT^{+}98]. Honour [Ano07, ACHS04c]. horticultural [Nel67]. horticulture [Nel66a]. Horton [Nel79c]. human [LN00b]. hybrid [CS05, LY12, RWJ13, SGK16, SEABZ24]. hybridized [CS03, LL12]. hyperbolic [Lóc13].

I/II [Lar97]. ICA [Sel15]. Identification [Nel63b]. IFIP [Fos79, GH92]. II [Lar97, Nel65b]. IK [YCH24]. image [LL12]. implementation [SS04]. Implementing [GH12]. importance [Nel98c]. important [Nel93b, NR97]. Improved [Meh20, ZZP^{+}22]. Improvement [KiANS22, LN98, LNP11, Nel95a, LN99b]. Improving [SEABZ24]. Incomplete [TFB21]. Independence [Ait50, Ait59, Nel51]. Index [CCG12]. inequality [AT18]. Inference [Paw04, Wak04, LN09a, LN09b]. Informatics [DLPT80, Did84]. information [Nel68d, NMD^{+}98]. Informatique [Did84]. Informative [NMS98]. ingredients [Pay04]. initialization [TOO23, Wes19]. initio} [PP23]. innovative [Pay04]. Input [NC71]. Input/Output [NC71]. Instances [LN82]. Institut [DLPT80, Did84]. integer [Bre13]. Intelligence [Nel87a, Phe87, Str88]. Intelligent [Nel76a]. Interactions [Phe87]. Interactive [Nel75a, BN78, Nel84a]. Interface [Bil85, BS86, SMNS01]. International [Ano75, DLPT80, Did84, Hau86]. Interpretation [Nel85c, Nel54a, Nel63c]. Introduction [CCG12, Nel75c]. Introductory [Nel74c]. Invariance [Nel68e]. Inverse [Nel66b, Nel68f, YCH24, NM98]. inversion [KiANS22]. Invited [Ano75]. isotropic [HN55]. iterations [SS01]. Iterative [Nel80c, KiANS22].

J [Alt80, Ano07, Boo07, Bur91, Che90, Col79, Dal72, Lac90, Nel64a, Nel72b, Nel79d, Nel80a, Nel87b, Nel94c, Pla84, Pre84, Rob07, Sch92, Wat86, Wil84]. J. [Ait95, Bai04, Gow95, RTW95, Sea95]. Jacobian [YCH24]. JANM [YCH24]. JANM-IK [YCH24]. Jarvis [Nel63d]. John [ACHS04c, Ano07, Dry10, Gar71, Gho08, Inn70, K.71, Nel63a, Nel78a, Van95, ACHS04b, Ano11, MGP11, Sen03, Sen04, Sen19]. Joint [Nel92b, NLB^{+}98, LN99b, SGK16]. Joules [NZE^{+}15]. journals [Alt00, Nel01b]. Jr [Nel80a]. Julian [Ano11].

Karlsruhe [GH92]. Kempthorne [Nel72c]. Kennedy [Nel80a]. kind [Nel77c]. Kinds [Nel62b]. Kinematics [YCH24]. Kinetic [NRCC91]. knowledge [Nel91, WON88]. knowledge-based [Nel91, WON88]. Konferenz [FU88].

L [Nel75b, Nel79c, Nel82b, Nel98b]. laboratory [Nel59, NBCS66]. Lancashire [Jef72]. Languages [Nel71b, NG71, Nel84a, Pay04]. Large [Rip04, eSAdSM18, Nel98d]. Leading [JK97]. least [Nel74b, Nel80c]. Lee [Boo07, Dry10, Gho08, Rob07]. Leiden [CH78]. Lellouch [Nel64a]. Lemna [GN64]. Leroy [Nel72c]. Less [eSAdSM18]. Letter [Nel98b, Nel01b, NLSS08]. Letters [BLN^{+}81, NS74, NKA92, NR97, Nel98c, YYB^{+}76]. lettuce [NM56a, NM56b]. level [GH92, KBČ20]. Lévy [Baş26]. library [LIN88]. light [NM56a]. Likelihood [Lee04, LN05, LN09a, NP87, NL92, Paw04, Des97a, GT89, LN00b, LNN07, Les97, LR97, Nel85a, NKA92, Nel94c, Nel97c, NL97, Nel00c, Wir97, LN09b, Rob07, Boo07, Dry10]. Likelihood-based [Lee04]. Likhachev [IBV^{+}22]. lily [NT02]. Limited [Aki77]. Linear [Boo07, Bur91, Che90, CST04, Gho08, Lac90, LNP06, LNP17, MN83, MN89, NW72, Nel75a, Nel77a, Nel79c, Nel82a, NB82a, Nel82c, Nel82d, Nel87a, NRCC91, Pla84, Pre84, Rob07, Sch92, Sen04, Wak04, Wat86, Wil84, Ait95, BN78, Bre04, Dry10, Gow95, LN96, LN98, LN01, LN02, LN06a, Nel74b, Nel77d, NL91, NGN92, NL94, Nel94d, Nel95a, Nel95b, Nel95c, NV97, Nel98d, NL04a, NL04b, RTW95, Sea95, Van95]. Lines [Nel90a]. localization [KNCS21]. Log [Nel74b]. logic [LTJ95]. Logistic [Ano61, Nel62a, Nel61a]. Longford [Nel94b]. longitudinal [BC97, Nel98b]. lot [SGK16]. Low [LRWW98, LRWW99, Lóc13]. lucerne [Nel63d].

M [Ano07, Nel78a, Nel87b, Nel94c]. MacLagan [Nel75d]. magnetic [NM98]. Main [Nel68a]. making [SMNS99]. Makov [Nel94c]. Manual [BN78, Alt80]. Manuals [Nel78c]. manurial [ANB64]. many [Nel80c]. March [Bil85, BS86, Jef72]. marginal [LN04]. marginality [NR97, Nel98c, Nel00a]. Martin [Ano07]. material [PP23]. mathematical [ANB64, Uli58, Jef72]. Mathematics [Nel72d]. Matrices [TFB21]. Matrix [Gol69]. Maxwell [Mas09b]. McCullagh [Bur91, Che90, Lac90, Pla84, Pre84, Sch92, Wil84, Wat86]. Mead} [PP23, Aki77, Arn93, AT18, Baş26, Bre04, Bre13, BPT06, Cha12, CS03, CS05, CCG12, CGVP15, DCZ^{+}19, DDM^{+}16, FBP19, FZ06, Gal22, Gal24, GH12, eSAdSM18, HNX03, HN06, IBV^{+}22, KTL19, Kel99, KiANS22, KBČ20, KNCS21, LRWW98, LRWW99, LPW12, Lar97, LTJ95, LL12, LY12, Lóc13, MM26, Mas06, Mas07, Mas09a, Mas09b, McK98, McK99, Meh20, MJ10, NZE^{+}15, NT02, ON75, OT09, Per24, Per26, PCB02, Pri21, PB26, RWJ13, SGK16, Sel15, SS01, SS04, SEABZ24, TOO23, TFB21, Tom95, TA13, WYWW15, Wes19, DW87, Wri12, WSKF13, XD14, YCH24, ZHZ^{+}07, ZZP^{+}22]. Mean [NLB^{+}98, Nel92b, LN99b]. means [KTL19, LL12, NZE^{+}15]. measurement [ZHZ^{+}07]. Measuring [NLW85]. mechanical [DCZ^{+}19]. medical [Alt00, Nel01b]. Medicine [Nel98b, Nel01b]. Meeting [BS^{+}97]. Memorian [NL95]. Mesh [AT18]. Mesh-based [AT18]. metamaterial [DCZ^{+}19]. Method [Aki77, CCG12, LRWW98, McK98, Bre13, Cha12, CS05, DCZ^{+}19, Gal22, eSAdSM18, HNX03, HN06, IBV^{+}22, KTL19, KiANS22, LRWW99, Lar97, Lóc13, Mas06, Mas07, McK99, NZE^{+}15, NM65a, NM65b, OT09, Pri21, RWJ13, Sel15, TOO23, WYWW15, Wri12, XD14, ZHZ^{+}07, ZZP^{+}22]. Methods [ACHS04c, BJM14, NBCS66, Nel78a, Nel80b, Phe87, Mas09a, Ano07]. Microcomputers [DW87]. Milton [Dal72, Gar71, Inn70, K.71]. Minimisation [Aki77]. Minimization [CCG12, ON75, TFB21, NM65a, NM65b, Sel15]. minor [GN64]. Missing [Nel54b]. Mixed [CST04, NLSS08, Bre13, Nel97b]. mixed-integer [Bre13]. mixed-model [Nel97b]. Mixed-Models [NLSS08]. Model [Cox04, CST04, Nel68e, Nel79c, ANB64, IBV^{+}22, Nel97b, NM98, SMNS99, SNB^{+}01, SMNS01, Nel77a]. Model-Building [Nel68e]. model-fitting [SMNS01]. Modeling [NLB^{+}98]. modellers [NMT^{+}98]. Modelling [GFNF82, Nel75a, Nel87a, Nel92c, NMT^{+}98, NL01, Paw04, Wak04, BN78, Des97a, LN00b, LN00c, Nel92b, Nel97c, LN99b, Les97, LR97, Wir97]. Models [Ano07, Boo07, Gho08, Lee04, LNP06, LNP17, MN83, MN89, NW72, Nel82a, NB82a, NF84, Nel85b, Nel86a, NRCC91, Nel92d, Nel98e, NLSS08, Rip04, Rob07, Sen04, ACHS04c, Ait95, BC97, Dry10, Gow95, Jef72, LN96, LN98, LN01, LN02, LN04, LN05, LN06a, LN06b, LN09a, LN09b, Nel53, Nel74b, Nel77d, Nel82d, Nel83a, NL91, NGN92, NL94, Nel94d, Nel95a, Nel95b, Nel95c, NV97, Nel98d, Nel98b, NMD^{+}98, NL04a, NL04b, RTW95, Sea95, Van95, Nel82c, Bur91, Che90, Lac90, Nel94b, Pla84, Pre84, Sch92, Wat86, Wil84]. Modified [NM98, Arn93, DDM^{+}16, Lar97, Pri21]. Monitoring [PV99]. Morgan [Nel87b]. most [Nel93b]. MR1229290 [Nel94c]. MR2757427 [LN09b]. MR2757428 [LN09b]. MR2757429 [LN09b]. MR2757430 [LN09b]. MR3363409 [NM65a]. muddle [Nel97b]. Multi [Nel66b, Nel68a, Nel77b, XD14]. Multi-Dimensional [Nel77b]. Multi-direction-based [XD14]. Multi-Factor [Nel66b]. Multi-Way [Nel68a]. multidimensional [Mas07]. multiminima [CS03, CS05]. Multiresponse [Nel72b]. My [Nel77c].

N [Nel72b, Nel94b]. NAG [LIN88]. nanoparticle [ZHZ^{+}07]. National [Did84, Nel59]. Nearly [Nel90a]. need [Nel84a]. negative [Nel54a]. neighbourhood [DDM^{+}16]. Nelder [ACHS04c, Ait95, Ano07, Ano11, Boo07, Dry10, Gho08, Gow95, PP23, Pre84, Rob07, RTW95, Sea95, Sen19, Wat86, ACHS04b, Aki77, Alt80, Arn93, AT18, Bai04, Baş26, Bre04, Bre13, BPT06, Bur91, Cha12, Che90, CS03, CS05, CCG12, Col79, CGVP15, Dal72, DCZ^{+}19, DDM^{+}16, Dvo57, FBP19, FZ06, Gal22, Gal24, GH12, Gar71, eSAdSM18, Ham52, HNX03, HN06, Inn70, IBV^{+}22, K.71, KTL19, Kel99, KiANS22, KBČ20, KNCS21, Lac90, LRWW98, LRWW99, LPW12, Lar97, LTJ95, LL12, LY12, Lóc13, MM26, Mas06, Mas07, Mas09a, Mas09b, MGP11, McK98, McK99, Meh20, MJ10, NZE^{+}15, NT02, ON75, OT09, Per24, Per26, Pla84, PCB02, Pri21, PB26, RWJ13, SGK16, Sch92, Sel15, Sen03, Sen04, SS01, SS04, SEABZ24, TOO23, TFB21, Tom95, TA13, Uli58]. Nelder [Van95, WYWW15, Wes19, Wil84, DW87, Wri12, WSKF13, XD14, YCH24, ZHZ^{+}07, ZZP^{+}22]. Netherlands [BvdH00]. networks [KNCS21]. Newsletter [Nel88c]. Newtonian [NZE^{+}15]. next [Nel76a, Nel93b]. Niall [Ano07]. nitrogen [GFNF82]. no [Nel94c, Nel00d]. nodes [PP23, PP23]. Non [Wak04, DDM^{+}16, LN00a, LN00c, NL01]. non-differentiable [DDM^{+}16]. Non-Linear [Wak04]. non-normal [LN00a, LN00c, NL01]. nonexistence [Str88]. Nonlinear [CCG12, Per24, Per26, Bre13, RWJ13]. nonorthogonal [Nel82d]. Nonsmooth [Sel15]. Nonstationary [McK98, McK99]. Normal [Ait50, Ait59, LN00a, LN00c, NL01]. North [Nel87b]. Note [CGVP15, Nel54b, Nel64b, Nel51, Nel61b, Nel62a]. Notes [Nel56b]. null [Nel65a]. Numerical [KTL19, Fos79].

Obituaries [Ano11]. Obituary [Nel90b]. occasions [Nel80c]. occurring [Nel52]. October [DLPT80, Did84, Sen19]. official [Nel83a]. Old [Gal24]. One [Nel77b, Nel00e]. Open [Gal24]. Optical [CCG12]. Optimal [Baş26, Des97a, Nel97c, Lar97, PB26, Les97, LR97, Wir97]. Optimization [PV99, DW87, AT18, Bre04, Bre13, Cha12, CS03, CS05, DDM^{+}16, FBP19, FZ06, eSAdSM18, LY12, Mas07, Mas09a, Per24, Per26, Pri21, SEABZ24, TOO23, Tom95, ZZP^{+}22]. optimize [IBV^{+}22, SGK16]. optimizer [Baş26]. Optimizing [Arn93, LTJ95]. ordinary [Mas06]. organised [Did84]. organism [Nel61b]. organized [DLPT80, Hau86]. Oriented [Nel79b]. origin [Nel92d]. Ornithology [Nel87b]. orthogonal [Nel65a, Nel65b]. Oscar [Nel72c]. other [LIN88, NABS60, Nel88a, Wri12]. outliers [Nel00d]. Output [NC71]. overcurrent [LY12]. overdispersion [LN00c]. oxygen [GGN64].

P [Bur91, Che90, Lac90, Nel79b, Nel87b, Nel98b, Pla84, Sch92, Wat86, Wil84]. P. [Pre84]. Packages [Nel74d, Nel74c, Nel92e, Nel00b]. pages [Nel82b]. Pairwise [TFB21]. Panel [NAM^{+}92]. Paper [Nel88c, Car97, Les97, LR97, Nel89, Nel97c, Wir97]. Papers [Bai04, Hau86, Nel83b, Ano75]. Parabolic [CCG12]. Parabolic-Index [CCG12]. Parallel [Nel90a]. parallelisation [KBČ20]. Parameter [CGVP15, NAM^{+}92, PB26]. parameters [GH12, Meh20, NGN92, TA13]. parsimony [LN03b]. particle [LY12]. Partitioned [MM26]. Party [Nel75a]. past [Nel93a]. Patterns [Cal71, Nel61b]. Pawitan [Dry10, Boo07, Gho08, Rob07]. Performance [Fos79]. Permutation [BJM14]. Personal [Nel84b]. Personalities [JK97]. Perspectives [CST04, Paw04]. perturbed [FBP19]. Phase [Lar97]. PIDD [Baş26]. Plackett [Nel75b]. planning [Nel56b]. Plant [BN60, Nel62c]. Plot [Nel54b]. Plots [Nel90a]. Point [McK98, KTL19, McK99, NZE^{+}15, Nel72e]. points [Meh20]. Poisson [NF84]. Polynomials [Nel66b, Nel68f, HNX03, NM98]. population [BN60]. Position [Nel84b]. Potential [Nel84b]. powerful [Nel95a]. Practical [TOO23]. Precision [Aki77]. Prediction [LN82]. Preface [ACHS04a]. Present [JK97, Nel84b, Nel93a]. Presentation [Nel83b]. Price [Nel79a]. Principle [Nel98e]. Principles [Bai04].

Probability [Nel72c]. problem [Dvo57, GH92, KTL19, MM26, Mas07, NMT^{+}98, SGK16]. problem-solving [MM26]. problems [Bre13, LNN07, Mas09a, OT09, RWJ13, TOO23]. Procedure [ON75, FZ06, Per24, Per26]. Proceedings [BvdH00, BFS74, CH78, DLPT80, Ano75, BS^{+}97, Bil85, BS86, Did84, Fos79, GH92]. Process [PV99, HN55, LN99c]. Professor [Ano07, ACHS04c]. Professors [Nel89]. Program [Nel75a, AGL^{+}70]. programmes [Nel76a]. Programming [GH92, NC67, NC71, RWJ13]. progress [NB85]. Proper [Wes19]. Properties [LRWW98, LRWW99]. prospects [NB85]. Pseudo [NKA92, Nel00c]. Pseudo-likelihood [NKA92, Nel00c]. pseudolikelihood [NL92]. Pump [NGMS94, Nel94c]. Pump-Failure [NGMS94, Nel94c]. purpose [NG71]. PY [PP23]. PY-Nodes [PP23]. Python [PP23].

Quadratic [Ait50, Ait59, Nel51]. Qualitative [Nel85b]. qualités [Nel80d]. qualities [Nel80d]. quality [LN98, LNP11, Nel95a, LN99b, Nel05]. quality-improvement [LN98, Nel95a, LN99b]. Quantal [Nel68f]. Quantitative [Nel62c, Nel72b, Nel60]. quantum [ZZP^{+}22]. Quasi [Nel85a, NP87, Nel00c, Des97a, GT89, LN00b, Les97, LR97, NKA92, NL92, Nel97c, NL97, Wir97]. Quasi-Likelihood [NP87, Nel85a, Nel00c, Des97a, GT89, LN00b, Les97, LR97, NKA92, NL92, Nel97c, NL97, Wir97]. quasilikelihood [LN99a]. Questions [Gal24, NGL^{+}94].

R [Alt80, Col79, Dal72, Nel64a, Nel72b, Nel75b, Nel79c]. Raktoe [Nel82b]. Ralston [Nel80b]. Random [Boo07, Gho08, LNP06, LNP17, Rob07, Dry10, LN01, LN05, LN06b, Nel94b]. random-effect [LN01, LN05, LN06b]. randomized [Nel65a, Nel65b]. rapid [SEABZ24]. Rarities [Nel62d, Nel04b]. Rates [NF84]. ratio [LN00b]. Re [NGMS94, Nel94c]. Re-Analysis [NGMS94, Nel94c]. Reaction [NGMS94]. Reader [NGMS94]. real [Baş26, MM26]. real-coded [MM26]. Recherche [DLPT80, Did84, Nel72a, NPT^{+}74, Nel75e]. recording [NBCS66]. reference [Nel80d, Nel80d]. reformulation [Nel77d]. Regenerative [Aki77]. Region [Hau86, Lar97]. Regression [Nel68e, Nel68f, Nel85c, Wak04, BC97, Nel98b, Nel79a]. regulatory [LN99c]. Rejoinder [Des97b, LN09b, Nel95b]. Relating [CGVP15]. relations [Nel63d]. relationship [LN00b]. relays [LY12]. Release [Alt80, BN78, Col79]. Remarkes [Cox04]. remarks [Nel74c, Str88]. Remediation [Kel99]. REML [CST04, LN03a]. Replicates [Nel63b]. Research [Nel59, Nel72a, NPT^{+}74, Nel75e, Nel67, Nel93b]. Research/Formation [Nel72a, NPT^{+}74, Nel75e]. Residual [Nel90a]. respiration [GGN64]. Response [Nel66b, Nel68f, Nel77b, NGMS94, NMD^{+}98, Nel98e, Nel00e, NLSS08, FZ06, GFNF82, Nel63c, NM98, Nel00a]. Response-Surface [Nel98e, NMD^{+}98, NM98, Nel00a]. restrained [BPT06]. Restricted [LPW12, Nel69b]. Results [Gal24]. Review [Ano07, Boo07, Bur91, Col79, Dal72, Dry10, Gar71, Gho08, Inn70, K.71, Lac90, Nel56a, Nel63a, Nel64a, Nel68c, Nel71a, Nel72b, Nel72c, Nel75b, Nel77a, Nel78a, Nel78c, Nel79c, Nel79a, Nel79b, Nel80b, Nel80a, Nel87b, Nel94b, Pla84, Pre84, Rob07, Sch92, Wat86, Wil84, Nel82b]. Reviews [Alt80, Che90, Nel79d]. RFIC [WSKF13]. Riemannian [Sel15]. Robert [Nel75d]. Robust [NL04b]. robustness [LN99a]. Role [Nel88b, Nel83a]. roots [HNX03]. Round [Nel00e]. Rouquette [Nel64a]. Roy [Gar71, Inn70, K.71, Nel72b]. rules [Nel98c].

S [Nel72b, Nel79a, Nel79b, Nel80b]. same [Nel00c]. Sampling [HN55, SNB^{+}01]. Sands [Jef72]. Scaling [Nel76b]. Scand [Nel94c]. scatter [SGK16]. scheme [KBČ20]. schemes [SNB^{+}01]. Schrödinger [Mas09b]. Schwartz [Nel64a]. Science [Nel86b, Nel99, Bil85, BS86, LN99c, Nel98a]. Sciences [FKJ^{+}89, JK97, Nel79c]. scientific [GH92]. se [Nel80d]. search [Cha12, CS05, DDM^{+}16, FZ06, NT02, Per24, Per26, SGK16, SS01, SS04, Wes19]. searching [PP23, WYWW15]. Second [DLPT80]. section [NT02]. see [Nel88a]. seepage [OT09]. segmentation [LL12]. Selected [Hau86]. Selecting [Rip04, IBV^{+}22]. Selection [Nel98e]. Self [Aki77]. Semi [CCG12]. sensor [KNCS21]. September [BS^{+}97, GH92]. servant [Nel80d]. Seventeenth [JK97]. sex [LN00b]. shape [GGN64]. shark [Baş26]. should [Nel88a]. signals [WYWW15]. Simple [Nel76b, Nel61b]. Simplex [Aki77, CCG12, Gal24, LRWW98, McK98, ON75, DW87, Arn93, Cha12, CS05, FBP19, FZ06, GH12, HNX03, HN06, LRWW99, Lar97, LTJ95, LL12, Lóc13, McK99, NM65a, NM65b, OT09, PP23, Per24, Per26, PB26, RWJ13, Tom95, TA13, Wes19, Wri12, ZHZ^{+}07, ZZP^{+}22]. simplified [eSAdSM18]. simulated [TA13]. Simulation [ZHZ^{+}07, Cha12]. Singular [Nel85c, Mas07]. Singular-Value [Nel85c]. Sixteenth [Bil85]. Sixty [Gal24]. size [ZHZ^{+}07]. sizing [SGK16]. smart [KNCS21]. Smith [Nel94c]. Social [Nel79c]. Society [Hau86, Jef72]. Software [FU88, Nel88c, Fos79, Nel78b, NB85]. soil [GGN64]. Solution [Mas09b, KTL19, Mas06, Mas07]. solutions [LNN07]. solve [SEABZ24]. solving [GH92, MM26, RWJ13]. Some [Cal71, Cox04, Nel52, Nel84b, Alt00, Nel61b, NL92, Nel01b, Van95]. souhaitables [Nel80d]. source [TA13]. Spacing [Nel62b, Meh20, NM56a, NM56b]. special [Nel84a]. Splicing [Nel94a]. SPSS [Nel78c]. squares [Nel74b, Nel80c]. Srivastava [Nel72b]. stackloss [Nel00d]. stage [Nel76a]. Stagnation [Kel99, NZE^{+}15]. Standardization [LN82]. standardized [NGN92]. Stanley [Nel63a]. Static [KNCS21]. statics [KTL19]. Station [Nel59]. Statist [Nel94c]. Statistical [Ait50, Ait59, BJM14, FKJ^{+}89, Gol69, JK97, MN69, MN14, Nel53, NC67, Nel69b, NC71, Nel71b, NG71, Nel72a, Nel74c, NPT^{+}74, Nel74d, Nel75a, Nel75e, Nel76b, Nel84b, Nel85b, NB85, Nel88c, Nel92e, NGL^{+}94, Nel99, PV99, Paw04, Phe87, Des97a, Les97, LR97, LN99c, Nel51, Nel62d, Nel69a, Nel74a, NP75, Nel76a, Nel78b, Nel79e, Nel80d, Nel88a, Nel92a, Nel93a, Nel93b, Nel97c, Nel98a, Nel04b, Wir97, WON88, CJM^{+}62, Nel62c, Dal72, Gar71, Inn70, K.71, Nel79b, Nel80b, Nel80a]. statistician [Nel72e].

Statisticians [BS^{+}97, Nel84a]. Statistics [Alt00, Ano07, BvdH00, BFS74, BT91, CH78, Hau86, LN99c, Nel72c, Nel86b, Nel88b, Nel99, ACHS04c, Ait95, AGL^{+}70, BS^{+}97, Bil85, BS86, Gow95, Nel77c, Nel82c, Nel83a, Nel92c, Nel94d, Nel95b, Nel95c, Nel98a, Nel00b, RTW95, Sea95, Str88, Uli58, Van95, Nel98b, Nel01b, Nel71a, Nel87b]. Statistik [FU88]. Statistik-Software [FU88]. Statistiques [Nel72a, NPT^{+}74, Nel75e, Nel80d]. steepest [Arn93]. Step [CCG12]. Stephens [Ano07]. Stochastic [Cha12, FZ06, Arn93, Tom95]. Storage [Nel69b]. Streitberg [Nel88c]. stretching [NZE^{+}15]. Strong [Nel98e]. structure [Nel65a, Nel65b]. structured [LN01]. structures [NM56a, Nel69a, NP75, Pay04]. studies [BC97, Nel98b]. study [NABS60]. subject [Lar97]. subspace [KiANS22]. Sufficient [Kel99]. suitable [WSKF13]. Summary [Nel72e]. Surface [Nel98e, FZ06, NZE^{+}15, NM98, NMD^{+}98, Nel00a]. surfaces [Arn93, Nel63c]. SVD [Mas07]. swarm [LY12]. Symposium [BvdH00, Bil85, BS86, DLPT80, Did84, CH78, Jef72]. Synthesis [Nel05, LN01, Nel82c]. System [Alt80, BN78, Col79, Nel87a, Nel74a, Nel88a]. Systematic [Nel62b]. systèmes [Nel80d]. Systems [Hau86, Nel74d, Nel88b, KTL19, NG71, Nel80d, Str88].

T [Nel56a, Nel82b, Nel87b, Nel94b]. Table [Nel68b, Nel68a, Nel77b]. Tables [Nel00e, Nel74b]. tabu [CS05]. Tabulated [Nel79d]. Taguchi [NAM^{+}92, NL91]. Taguchi-type [NL91]. TC [Fos79]. TC2 [GH92]. TC2/WG [GH92]. teaching [Nel00b]. Technical [CGVP15]. Technology [Nel86b]. ten [Nel93b]. tenability [Van95]. tensors [Mas07]. Terms [Nel98e]. their [CGVP15]. theoretical [Nel61b]. Theory [Nel56a, NV97, Nel77a]. thérapeutiques [Nel64a]. therapy [Lar97]. There [Nel00d]. thing [Nel00c]. Third [Did84]. Thompson [Nel89]. three [Arn93, KBČ20]. three-dimensional [Arn93]. three-level [KBČ20]. tool [Nel95a]. tools [SMNS99]. Topics [Nel71c]. toxicity [Lar97]. Training [Nel72a, NPT^{+}74, Nel75e]. Treatment [Nel65b]. trends [Alt00, Nel01b]. trial [Lar97]. Triangular [Nel69b]. Two [LPW12, LN00c, Nel00e, KTL19]. two-point [KTL19]. type [NL91]. types [Nel60].

U [Nel94c]. ulcer [LN02]. ultrafine [ZHZ^{+}07]. unbalanced [Nel92e]. unconfined [OT09]. Unconstrained [CCG12, eSAdSM18, Tom95]. underwater [KNCS21]. Unified [Boo07, Rob07, Dry10]. Unmeasurable [NLW85]. unobservables [LN09a, LN09b]. Unstable [Mas06]. Upton [Nel79d]. use [ANB64, Lar97, Nel63c]. Useful [Nel66b, Nel82c]. User [Nel74d, Nel88a]. Using [Aki77, Kel99, KiANS22, KNCS21, LN02, LTJ95, Nel80d, Nel82a, NMD^{+}98, PP23, Sel15]. Utrecht [BvdH00].

Value [Nel85c, Mas07, Mas09a]. Values [CGVP15, Nel54b, NMD^{+}98, PB26]. variable [DDM^{+}16]. Variance [Nel71c, Nel87c, Nel51, Nel52, Nel54a, Nel60, Nel65a, Nel65b]. variant [Lóc13, NT02, PCB02]. Variates [Ait50, Ait59]. variation [CJM^{+}62, NABS60]. variational [Mas09a]. Vegetable [Nel59]. vehicle [Baş26]. Versailles [DLPT80, Did84]. versus [LN99c, TA13]. via [Baş26, Boo07, Dry10, FZ06, LN06b, Mas06, Mas07, Mas09b, NL04b, Rob07]. View [Nel94a, LN04, LN09a, LN09b, Nel72e]. Views [Nel84b]. visual [SMNS01]. Visualisation [SMNS99]. visualization [SNB^{+}01]. Volkov [IBV^{+}22].

W [Nel80a, Nel82b]. Walter [Nel56a]. wave [DCZ^{+}19]. Way [Nel68a]. ways [LIN88, LN00c]. Weak [Nel98e]. Weak-Heredity [Nel98e]. weather [ANB64]. Wedderburn [Nel75d]. Weighted [Nel68f, Nel80c]. WG [GH92]. white [Baş26]. Wiley [Nel82b]. Wilf [Nel80b]. William [Nel75d]. wissenschaftliche [FU88]. within [LIN88]. Working [Fos79, GH92, Hau86, Nel75a]. Works [ACHS04b]. Workshop [Hau86].

Yates [NL95, Nel97a]. yearly [NABS60]. Years [Gal24, Nel93b]. Yield [Nel63d, BN60]. Yield-density [Nel63d]. yields [NABS60]. York [Nel82b]. Youngjo [Gho08, Dry10]. Yudi [Dry10, Gho08].

References

Adams:2004:P

[ACHS04a] N. M. Adams, M. J. Crowder, D. J. Hand, and D. A. Stephens. Preface. In Adams et al. [ACHS04c], pages ix–x. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Adams:2004:WJN

[ACHS04b] N. M. Adams, M. J. Crowder, D. J. Hand, and D. A. Stephens. The works of John Nelder. In Adams et al. [ACHS04c], pages 235–244. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (ebook). LCCN QA276.A1 M48 2004.

Adams:2004:MMS

[ACHS04c] Niall Adams, M. J. (Martin J.) Crowder, D. J. (David J.) Hand, and David Stephens, editors. Methods and models in statistics: in honour of Professor John Nelder FRS. Imperial College Press, London, UK, 2004. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). xiii + 246 pp. LCCN QA276.A1 M48 2004.

Alvey:1970:GGS

[AGL^{+}70] N. G. Alvey, John C. Gower, F. Bruce Lauckner, P. K. Leech, John Ashworth Nelder, C. E. Rogers, Gavin Ross, H. R. Simpson, Robert William MacLagan Wedderburn, and Graham Neil Wilkinson. GenStat: a general statistics program. ????, ????, 1970.

Aitken:1950:SIQ

[Ait50] A. C. Aitken. On the statistical independence of quadratic forms in normal variates. Biometrika, 37(1/2):93–96, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332151. See note [Nel51] and

corrigenda [Ait59].

Aitken:1959:CSI

[Ait59] A. C. Aitken. Corrigenda: “On the Statistical Independence of Quadratic Forms in Normal Variates”. Biometrika, 46(1/2):279, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332856. See

[Ait59].

Aitkin:1995:CJN

[Ait95] Murray Aitkin. Comments on J. A. Nelder: ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):85–86, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00143934. See [Nel94d].

Akitt:1977:FMU

[Aki77] J. W. Akitt. Function minimisation using the Nelder and Mead simplex method with limited arithmetic precision: the self regenerative simplex. The Computer Journal, 20(1):84–85, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/20/1/84.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/84.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/85.tif.

Altman:1980:BRG

[Alt80] D. G. Altman. Book reviews: The GLIM System Release 3 Manual, by R. J. Baker and J. A. Nelder. Applied Statistics, 29(1): 98, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Altman:2000:SMJ

[Alt00] Douglas G. Altman. Statistics in medical journals: some recent trends. Statistics in Medicine, 19(23):3275–3289, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [Nel01b].

Austin:1964:UMM

[ANB64] R. B. Austin, J. A. Nelder, and G. Berry. The use of a mathematical model for the analysis of manurial and weather effects on the growth of carrots. Annals of Botany, 28(1):153–162, January 1964. CODEN ANBOA4. ISSN 0305-7364 (print), 1095-8290

(electronic). URL https://academic.oup.com/aob/article-
abstract/28/1/153/182917.

Anonymous:1961:CFG

[Ano61] Anonymous. Correction: The fitting of a generalization of the logistic curve. Biometrics, 17(3):509, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527846. See [Nel61a].

Anonymous:1975:IPP

[Ano75] Anonymous, editor. Invited papers: proceedings of the 9th International Biometric Conference, Boston, August 22–27, 1976, volume II. Biometric Society, Boston, MA, USA, 1975. ISBN ???? LCCN QH323.5 .I55 1976.

Anonymous:2007:BRBs

[Ano07] Anonymous. Book review: Methods and Models in Statistics, in Honour of Professor John Nelder, FRS by Niall M. Adams; Martin Crowder; D. J. Hand; Dave Stephens. Technometrics, 49(4): 501, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471414.

Anonymous:2011:OJA

[Ano11] Anonymous. Obituaries: John Ashworth Nelder; Julian Besag. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):499–504, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Arnholt:1993:OTD

[Arn93] Alan Todd Arnholt. Optimizing three-dimensional surfaces with stochastic error distributions: a comparison between the steepest ascent algorithm and the modified Nelder–Mead simplex algorithm. ProQuest LLC, Ann Arbor, MI, 1993. 160 pp.

URL http://gateway.proquest.com/openurl?url_ver=Z39.
88-2004&rft_val_fmt=info:ofi/fmt:kev:mtx:dissertation&
res_dat=xri:pqdiss&rft_dat=xri:pqdiss:9323103. Thesis

(Ph.D.)–University of Northern Colorado.

Audet:2018:MBN

[AT18] Charles Audet and Christophe Tribes. Mesh-based Nelder–Mead algorithm for inequality constrained optimization. Computational optimization and applications, 71(2):331–352, 2018. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Bailey:2004:PDE

[Bai04] R. A. Bailey. Principles of designed experiments in J. A. Nelder’s papers. In Adams et al. [ACHS04c], pages 171–194. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Basak:2026:ORP

[Baş26] Hasan Başak. Optimal real PIDD^{2} controller design for vehicle cruise control via Lévy flight-based white shark optimizer with Nelder–Mead algorithm. International Journal of Dynamics and Control, 14(1):Paper No. 36, 31, 2026.

Broemeling:1997:BAR

[BC97] Lyle D. Broemeling and Peyton Cook. A Bayesian analysis of regression models with continuous errors with application to longitudinal studies. Statistics in Medicine, 16(4):321–332, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [Nel98b].

Bruckmann:1974:CPC

[BFS74] Gerhart Bruckmann, Franz Ferschl, and Leopold Schmetterer, editors. COMPSTAT: Proceedings in Computational Statistics. Physica-Verlag, Vienna, Austria, 1974. ISBN 3-7908-0148-8. ISSN 0253-018X. LCCN QA276.4 .C18 1974.

Billard:1985:CSS

[Bil85] Lynne Billard, editor. Computer science and statistics: proceedings of the Sixteenth Symposium on the Interface, Atlanta, Georgia, March 1984. North-Holland, Amsterdam, The Netherlands, 1985. ISBN 0-444-87725-8 (paperback). LCCN QA276.4 .S95 1984.

Berry:2014:CPS

[BJM14] Kenneth J. Berry, Janis E. Johnston, and Paul W. Mielke, Jr. A Chronicle of Permutation Statistical Methods: 1920–2000, and Beyond. Springer International Publishing, Cham, Switzerland, 2014. ISBN 3-319-02743-3 (print), 3-319-02744-1 (e-book). xix + 517 + 51 pp. LCCN QA165 .B47 2014; QA165 .C57 2014.

Beale:1981:LE

[BLN^{+}81] E. M. L. Beale, P. W. Lane, J. A. Nelder, A. F. Bissell, and P. L. Goldsmith. Letters to the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):231–234, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/stable/2988053.

Bleasdale:1960:PPC

[BN60] J. K. A. Bleasdale and J. A. Nelder. Plant population and crop yield. Nature, 188(4747):342, October 22, 1960. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:

//www.nature.com/articles/188342a0.

Baker:1978:GSM

[BN78] R. J. Baker and John A. Nelder. The GLIM System Manual, Release 3: Generalized linear interactive modelling. Numerical Algorithms Group, Oxford, UK, 1978.

Booth:2007:BRG

[Boo07] James G. Booth. Book review: Generalized Linear Models with Random Effects: Unified Analysis via H-likelihood, by Y. Lee, J. A. Nelder, and Y. Pawitan. Biometrics, 63(4):1296–1297, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burmen:2006:GRN

[BPT06] Árpád Bűrmen, Janez Puhan, and Tadej Tuma. Grid restrained Nelder–Mead algorithm. Computational optimization and applications, 34(3):359–375, 2006. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Brea:2004:NMO

[Bre04] Ebert Brea. Nelder–Mead optimization under linear constraints. PhD thesis, University of Southampton, Southampton, UK,

2004. ???? pp. URL https://www.proquest.com/pqdtglobal/

docview/305035514. Thesis (Ph.D.)–University of Southampton (United Kingdom).

Brea:2013:ENM

[Bre13] Ebert Brea. An extension of the Nelder–Mead method to nonlinear mixed-integer optimization problems. Revista Internacional de Métodos Numéricos para Cálculo y Diseño en Ingeniería, 29(3): 163–174, 2013.

Boardman:1986:CSS

[BS86] Thomas J. Boardman and Irene M. Stefanski, editors. Computer science and statistics: proceedings of the 18th Symposium on the Interface, Fort Collins, Colorado, March 1986. American Statistical Association, Washington, DC, USA, 1986. LCCN QA276.4 .S95 1986.

Barra:1997:RDS

[BS^{+}97] Jean Renée Barra, Helmut Strasser, et al., editors. Recent developments in statistics: proceedings of the European Meeting of Statisticians, Grenoble 6–11 September 1976. North-Holland, Amsterdam, The Netherlands, 1997. ISBN 0-7204-0751-6. LCCN QA276.A1 E89 1976.

Buja:1991:CGS

[BT91] Andreas Buja and Paul A. Tukey, editors. Computing and Graphics in Statistics, volume 36 of IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97633-7. LCCN QA276.4 .C59 1991.

Burridge:1991:BRB

[Bur91] Jim Burridge. Book review: Generalized Linear Models, by P. McCullagh; J. A. Nelder. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):361–364, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/2983054.

Bethlehem:2000:CPC

[BvdH00] Jelke G. Bethlehem and Peter G. M. van der Heijden, editors. COMPSTAT Proceedings in Computational Statistics 14th Symposium held in Utrecht, The Netherlands, 2000. Physica-Verlag, Heidelberg, Germany, 2000. ISBN 3-7908-1326-5 (paperback), 3-642-57678-8 (e-book). LCCN QA276.4 .C667 2000.

Calinski:1971:SDP

[Cal71] T. Caliński. On some desirable patterns in block designs. Biometrics, 27(2):275–292, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528995.

Carroll:1997:CPD

[Car97] R. J. Carroll. Comment on the paper by Desmond. Journal of Statistical Planning and Inference, 60(1):104–106, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897900100. See [Des97a].

Choudhury:2012:SAF

[CCG12] Raja Roy Choudhury, Arundhati Roy Choudhury, and Mrinal Kanti Ghose. Semi analytical formulation of dispersion characteristics for a step and parabolic-index optical fiber: Introduction of Nelder–Mead simplex method for nonlinear unconstrained minimization. International Journal of Computer Applications, 49(??):38–44, July 2012. CODEN ???? ISSN 0975-

8887. URL https://www.ijcaonline.org/archives/volume49/
number22/7906-1292/.

Corominas:2015:TNR

[CGVP15] Albert Corominas, Alberto García-Villoria, and Rafael Pastor. Technical note: Relating to the parameter values given by Nelder and Mead in their algorithm. The Computer Journal, 58(1):157– 159, January 2015. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/58/1/157.

Corsten:1978:CPC

[CH78] Leo Caspar Antoon Corsten and J. Hermans, editors. Compstat 1978, Proceedings in Computational Statistics: 3rd symposium held in Leiden 1978. Physica-Verlag, Wien, Austria, 1978. ISBN 3-7908-0196-8 (paperback). ISSN 0253-018X. LCCN QA276.4 .C192.

Chang:2012:SNM

[Cha12] Kuo-Hao Chang. Stochastic Nelder–Mead simplex method—a new globally convergent direct search method for simulation optimization. European Journal of Operational Research, 220(3):684–694, 2012. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Cheek:1990:BRG

[Che90] P. J. Cheek. Book reviews: Generalized Linear Models, 2nd Edn, by P. McCullagh and J. A. Nelder. Applied Statistics, 39(3):385– 386, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cooke:1962:ECC

[CJM^{+}62] Patricia Cooke, R. Morley Jones, K. Mather, G. W. Bonsall, and J. A. Nelder. Estimating the components of continuous variation. I. Statistical. Heredity, 17(??):115–133, February 1, 1962. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL

https://www.nature.com/articles/hdy19626.

Clarke:2000:C

[CNM00] Geoff Clarke, J. A. Nelder, and Douglas C. Montgomery. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):107–110, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.jstor.

org/stable/2681260.

Collett:1979:BRG

[Col79] D. Collett. Book review: The GLIM System Release 3, by R. J. Baker, J. A. Nelder. Biometrics, 35(2):527–528, June 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530359.

Cox:2004:SRM

[Cox04] D. R. Cox. Some remarkes on model criticism. In Adams et al. [ACHS04c], pages 13–22. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Chelouah:2003:GNM

[CS03] Rachid Chelouah and Patrick Siarry. Genetic and Nelder–Mead algorithms hybridized for a more accurate global optimization of continuous multiminima functions. European Journal of Operational Research, 148(2):335–348, 2003. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Chelouah:2005:HMC

[CS05] Rachid Chelouah and Patrick Siarry. A hybrid method combining continuous tabu search and Nelder–Mead simplex algorithms for the global optimization of multiminima functions. European Journal of Operational Research, 161(3):636–654, 2005. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Cullis:2004:PAR

[CST04] B. R. Cullis, A. B. Smith, and R. Thompson. Perspectives of ANOVA, REML and a general linear mixed model. In Adams et al. [ACHS04c], pages 53–94. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Dalenius:1972:BRB

[Dal72] Tore Dalenius. Book review: Statistical Computation by R. C. Milton; J. A. Nelder. International Statistical Review = Revue Internationale de Statistique, 40(2):254–256, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402775.

Dong:2019:AWB

[DCZ^{+}19] Jiaqi Dong, Weiqi Chen, Zhe Zeng, Qing-Hua Qin, and Yi Xiao. Analysis of wave band gaps in mechanical metamaterial based on Nelder–Mead method. Engineering Analysis with Boundary Elements, 103:109–115, 2019. ISSN 0955-7997 (print), 1873-197X (electronic).

Drazic:2016:CVN

[DDM^{+}16] Milan Dražić, Zorica Dražić, Nenad Mladenović, Dragan Urošević, and Qiu Hong Zhao. Continuous variable neighbourhood search with modified Nelder–Mead for non-differentiable optimization. IMA Journal of Management Mathematics, 27(1):75–88, 2016. CODEN IJMMCD. ISSN 1471-678X (print), 1471-6798 (electronic).

Desmond:1997:OEF

[Des97a] A. F. Desmond. Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):77–104, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896001231.

See discussion [Les97, Car97, Nel97c, Wir97, LR97] and rejoinder [Des97b].

Desmond:1997:R

[Des97b] A. F. Desmond. Rejoinder. Journal of Statistical Planning and Inference, 60(1):116–121, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900161. See [Des97a].

Diday:1984:DAI

[Did84] Edwin Diday, editor. Data analysis and informatics: proceedings of the Third International Symposium on Data Analysis and Informatics, organised by the Institut National de Recherche en Informatique et en Automatique, Versailles, October 4–7, 1983. North-Holland, Amsterdam, The Netherlands, 1984. ISBN 0-444-87555-7 (hardcover). LCCN QA278 .I56 1981.

Diday:1980:DAI

[DLPT80] Edwin Diday, Ludovic Lebart, J.-P. Pagés, and Richard Tomassone, editors. Data Analysis and Informatics: Proceedings of the Second International Symposium on Data Analysis and Informatics, organized by the Institut de Recherche d’Informatique et d’Automatique, Versailles, October 17–19, 1979. North-Holland, Amsterdam, The Netherlands, 1980. ISBN 0-444-86005-3 (hardcover). LCCN QA278 .I56 1979.

Drygas:2010:BRY

[Dry10] Hilmar Drygas. Book review: Youngjo Lee, John A. Nelder, Yudi Pawitan: Generalized linear models with random effects: unified analysis via H-likelihood. Statistical Papers, 51(4):1007– 1008, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00362-008-0175-7.

Dvoretzky:1957:PNH

[Dvo57] A. Dvoretzky. On a problem of Nelder and Hammersley. Bulletin of the Research Council of Israel. Section A. Mathematics, Physics and Chemistry, 6:115–118, 1957. ISSN 0366-2829.

Wouk:1987:OMN

[DW87] John E. Dennis, Jr. and Daniel J. Woods. Optimization on microcomputers: the Nelder–Mead simplex algorithm. In Arthur Wouk, editor, New Computing Environments: Microcomputers in Large-Scale Computing, pages 116–122. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1987.

Goncalves-e-Silva:2018:LMS

[eSAdSM18] Kayo Gonçalves e Silva, Daniel Aloise, Samuel Xavier de Souza, and Nenad Mladenović. Less is more: simplified Nelder–Mead method for large unconstrained optimization. Yugoslav Journal of Operations Research, 28(2):153–169, 2018. ISSN 0354-0243 (print), 1820-743X (electronic).

Fajfar:2019:NMS

[FBP19] Iztok Fajfar, Árpád Bűrmen, and Janez Puhan. The Nelder–Mead simplex algorithm with perturbed centroid for high-dimensional function optimization. Optimization Letters, 13(5):1011–1025, 2019. ISSN 1862-4472 (print), 1862-4480 (electronic).

Ferber:1982:ESS

[FKJ^{+}89] Robert Ferber, Samuel Kotz, Norman Lloyd Johnson, Campbell Read, et al., editors. Encyclopaedia of Statistical Sciences. Wiley, New York, NY, USA, 1982–1989. ISBN 0-471-05546-8 (vol. 1), 0-471-05547-6 (vol. 2), 0-471-05549-2 (vol. 3), 0-471-05551-4 (vol. 4), 0-471-05552-2 (vol. 5), 0-471-05553-0 (vol. 6), 0-471-05555-7 (vol. 7), 0-471-05556-5 (vol. 8), 0-471-81274-3 (vol. 10), 0-471-85474-3 (vol. 9). ???? pp. LCCN QA276.14 .E53 1981.

Fosdick:1979:PEN

[Fos79] Lloyd Dudley Fosdick, editor. Performance evaluation of numerical software: proceedings of the IFIP TC 2.5 Working Conference on Performance Evaluation of Numerical Software. Baden, Austria, 11–15 December 1978. North-Holland, Amsterdam, The Netherlands, 1979. ISBN 0-444-85330-8 (hardcover). LCCN QA297 .I18 1978.

Faulbaum:1988:FSS

[FU88] Frank Faulbaum and H.-M. (Hans-Martin) Uehlinger, editors. Fortschritte der Statistik-Software 1. 4. Konferenz über die wissenschaftliche Anwendung von Statistik-Software, Heidelberg, 1987, volume 1 of Eine ZUMA-Publikation. Gustav Fischer Verlag, Stuttgart, West Germany and New York, NY, USA, 1988. ISBN 3-437-50320-0. LCCN QA276.4 .K66 1987.

Fan:2006:SRS

[FZ06] Shu-Kai S. Fan and Erwie Zahara. Stochastic response surface optimization via an enhanced Nelder–Mead simplex search procedure. Engineering Optimization, 38(1):15–36, 2006. CODEN EGOPAX. ISSN 0305-215X.

Galantai:2022:CNM

[Gal22] Aurél Galántai. Convergence of the Nelder–Mead method. Numerical Algorithms, 90(3):1043–1072, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01221-7.

Galantai:2024:NMS

[Gal24] Aurél Galántai. The Nelder–Mead simplex algorithm is sixty years old: New convergence results and open questions. Algorithms (Basel), 17(11):1–21, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-

4893/17/11/523. See [NM65b].

Garside:1971:BRB

[Gar71] M. J. Garside. Book review: Statistical Computation, by Roy C. Milton; John A. Nelder. Journal of the Royal Statistical Society. Series A (General), 134(4):690–691, ???? 1971. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.jstor.org/stable/2343678.

Greenwood:1982:MCR

[GFNF82] Duncan Joseph Greenwood, M. J. S. Floate, J. A. Nelder, and M. J. Frissel. Modelling of crop response to nitrogen fertilizer. Philosophical Transactions of the Royal Society of London Series B, Biological Sciences, 296(1082):351–362, January 1982. CODEN PTRBAE. ISSN 0962-8436 (print), 1471-2970 (electronic).

Greenwood:1964:ESO

[GGN64] D. J. Greenwood, D. Goodman, and J. A. Nelder. Effect of shape on oxygen diffusion and aerobic respiration in soil aggregates. Journal of the Science of Food and Agriculture, 15(11):781–790, November 1964. CODEN JSFAAE. ISSN 0022-5142 (print), 1097-

0010 (electronic). URL https://scijournals.onlinelibrary.
wiley.com/doi/abs/10.1002/jsfa.2740151107.

Gaffney:1992:PEH

[GH92] P. W. Gaffney and Elias N. Houstis, editors. Programming environments for high-level scientific problem solving: proceedings of the IFIP TC2/WG 2.5 Working Conference on Programming Environments for High-level Scientific Problem Solving, Karlsruhe, Germany, 23–27 September, 1991, volume 2 of IFIP transactions. A, Computer science and Technology. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-89176-5 (paperback). ISSN 0926-5473. LCCN QA76.6 .I1782 1991. URL http://www.

gbv.de/dms/bowker/toc/9780444891761.pdf;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0746.00068.

Gao:2012:INM

[GH12] Fuchang Gao and Lixing Han. Implementing the Nelder–Mead simplex algorithm with adaptive parameters. Computational optimization and applications, 51(1):259–277, 2012. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Ghosh:2008:BRBb

[Gho08] Subir Ghosh. Book review: Generalized Linear Models with Random Effects by Youngjo Lee; John A. Nelder; Yudi Pawitan. Technometrics, 50(4):542–543, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471536.

Greenwood:1964:EDG

[GN64] D. J. Greenwood and J. A. Nelder. The effect of drugs on the growth of Lemna minor. Annals of Botany, 28(4):711–715, October 1964. CODEN ANBOA4. ISSN 0305-7364 (print), 1095-8290

(electronic). URL https://academic.oup.com/aob/article-
abstract/28/4/711/262098.

Golub:1969:MDS

[Gol69] G. H. Golub. Matrix decompositions and statistical computation. In R. C. Milton and J. A. Nelder, editors, Statistical Computation, pages 365–397. Academic Press, New York, NY, USA, 1969.

Gower:1995:CJN

[Gow95] John Gower. Comments on J. A. Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5 (2):91–92, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00143936. See [Nel94d].

Godambe:1989:EQL

[GT89] V. P. Godambe and M. E. Thompson. An extension of quasilikelihood estimation. Journal of Statistical Planning and Inference, 22(2):137–152, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375889901067.

See discussion [Nel89].

Hammersley:1952:CN

[Ham52] J. M. Hammersley. On a conjecture of Nelder. Compositio Mathematica, 10:241–244, 1952. CODEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic).

Haux:1986:ESS

[Hau86] Reinhold Haux. Expert Systems in Statistics: Selected Papers from a Workshop, organized by the Working Group “Computational Statistics” of the German Region of the International Biometric Society, December 6–7, 1985, Aachen, F.R.G. Gustav Fischer, Stuttgart, West Germany, 1986. ISBN 0-89574-235-7 (paperback), 3-437-40171-8 (paperback). 194 pp. LCCN QA276.4 .E925 1986.

Hammersley:1955:SIG

[HN55] J. M. Hammersley and J. A. Nelder. Sampling from an isotropic Gaussian process. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 51:652–662, 1955. CO-DEN PCPSA4. ISSN 0008-1981.

Han:2006:EDN

[HN06] Lixing Han and Michael Neumann. Effect of dimensionality on the Nelder–Mead simplex method. Optimization Methods and Software, 21(1):1–16, 2006. ISSN 1026-7670 (print), 1055-6788 (electronic).

Han:2003:RCP

[HNX03] Lixing Han, Michael Neumann, and Jianhong Xu. On the roots of certain polynomials arising from the analysis of the Nelder–Mead simplex method. Linear Algebra and its Applications, 363(C):109– 124, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print),

1873-1856 (electronic). URL http://www.elsevier.nl/gej-ng/
10/30/19/212/25/35/abstract.html. Special issue on nonnega-

tive matrices, M-matrices and their generalizations (Oberwolfach,

Ivanov:2022:ANM

[IBV^{+}22] A. M. Ivanov, F. S. Belyaev, A. E. Volkov, S. P. Belyaev, and N. N. Resnina. Application of the Nelder–Mead method to optimize the selecting of the Likhachev–Volkov model constants. Vestnik Sankt-Peterburgskogo Universiteta. Matematika. Mekhanika. Astronomiya, 9(67)(4):693–704, 2022.

Innis:1970:BRS

[Inn70] George Innis. Book review: Statistical Computation, by Roy C. Milton, John A. Nelder. Biometrics, 26(4):871, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528747.

Jeffers:1972:MME

[Jef72] John Norman Richard Jeffers, editor. Mathematical models in ecology: the 12th symposium of the British Ecological Society, Grangeover-Sands, Lancashire, 23–26 March 1971. Blackwell’s Scientific Publications, Oxford, UK, 1972. ISBN 0-632-08740-4. LCCN QH541.15M3 1972.

Johnson:1997:LPS

[JK97] Norman Lloyd Johnson and Samuel Kotz, editors. Leading Personalities in Statistical Sciences: from the Seventeenth Century to the Present. Wiley Series in Probability and Statistics. Wiley, New York, NY, USA, 1997. ISBN 0-471-16381-3 (paperback), 1-118-15071-6 (e-book), 1-118-15072-4, 1-283-28289-5. xxiii + 399 pp. LCCN QA276.156 .L43 1997.

K:1971:BRBe

[K.71] G. G. K. Book review: Statistical Computation by Roy C. Milton; John A. Nelder. Technometrics, 13(3):710–711, August 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267191.

Kriauziene:2020:TLP

[KBČ20] Rima Kriauzienė, Andrej Bugajev, and Raimondas Čiegis. A three-level parallelisation scheme and application to the Nelder– Mead algorithm. Mathematical Modelling and Analysis, 25(4):584– 607, 2020.

Kelley:1999:DRS2

[Kel99] C. T. Kelley. Detection and remediation of stagnation in the Nelder–Mead algorithm using a sufficient decrease condition. SIAM Journal on Optimization, 10(1):43–55, October/November 1999. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (elec-

tronic). URL http://epubs.siam.org/sam-bin/dbq/article/
31520.

Kitaoka:2022:INM

[KiANS22] Haru Kitaoka, Ken ichi Amano, Naoya Nishi, and Tetsuo Sakka. Improvement of the Nelder–Mead method using direct inversion in iterative subspace. Optimization and Engineering, 23(2):1033– 1055, 2022. ISSN 1389-4420 (print), 1573-2924 (electronic).

Kulandaivel:2021:SLU

[KNCS21] Madhumitha Kulandaivel, Arulanand Natarajan, Bharathi Priya Chandrasekaran, and Anandamurugan Selvaraj. Static localization for underwater acoustics sensor networks using Nelder–Mead algorithm for smart cities. Computational Intelligence, 37(4):1691– 1705, November 2021. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic).

Kalyukh:2019:NST

[KTL19] Yu. I. Kalyukh, A. N. Trofimchuk, and A. G. Lebed’. Numerical solution of a two-point statics problem for distributed extended systems by means of the Nelder–Mead method. Natsional’ naya Akademiya Nauk Ukrainy. Institut Kibernetiki im. V. M. Glushkova. Kibernetika i Sistemnyĭ Analiz, 55(4):109–118, 2019.

Lachenbruch:1990:BRG

[Lac90] P. A. Lachenbruch. Book review: Generalized Linear Models, by P. McCullagh, J. A. Nelder. Biometrics, 46(4):1231–1232, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532465.

Larsen:1997:EUN

[Lar97] Lois M. Larsen. An evaluation of the use of the Nelder–Mead modified simplex method to find the region of optimal dose therapy combinations subject to toxicity constraints in a Phase I/II clinical trial. Thesis (Ph.D.), University of Northern Colorado, Greeley, CO, USA, 1997. 196 pp. URL https://www.proquest.

com/pqdtglobal/docview/304369251.

Lee:2004:LBM

[Lee04] Y. Lee. Likelihood-based models beyond GLMs. In Adams et al. [ACHS04c], pages 195–214. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Lesperance:1997:DDP

[Les97] Mary L. Lesperance. Discussion on A. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):106– 108, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900112. See [Des97a].

Lane:1988:ANF

[LIN88] P. W. Lane, R. M. J. Iles, and John Ashworth Nelder. Accessing the NAG Fortran library from within GenStat, and other ways of extending GenStat. GenStat Newsletter, (20):41–50, 1988.

Li:2012:FMI

[LL12] Yan Ling Li and Ting Liu. Fuzzy means image segmentation based on hybridized Nelder–Mead simplex algorithm. Mathematics in Practice and Theory. Shuxue de Shijian yu Renshi, 42(14):218– 223, 2012. ISSN 1000-0984.

Lane:1982:ACS

[LN82] Peter W. Lane and John A. Nelder. Analysis of covariance and standardization as instances of prediction. Biometrics, 38(3):613– 621, September 1982. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/
2530043.

Lee:1996:HGL

[LN96] Y. Lee and J. A. Nelder. Hierarchical generalized linear models. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):619–678, 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://links.jstor.org/sici?sici=0035-9246(
1996)58:4<619:HGLM>2.0.CO;2-8&origin=MSN. With discus-

sion.

Lee:1998:GLM

[LN98] Y. Lee and J. A. Nelder. Generalized linear models for the analysis of quality-improvement experiments. The Canadian Journal of Statistics = La revue canadienne de statistique, 26(1):95–105, 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:1999:RQE

[LN99a] Y. Lee and John A. Nelder. The robustness of the quasilikelihood estimator. The Canadian Journal of Statistics = La revue canadienne de statistique, 27(2):321–327, 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Nelder:1999:JMM

[LN99b] Youngjo Lee and John A. Nelder. Joint modelling of the mean and dispersion for the analysis of quality-improvement experiments. In Park and Vining [PV99], chapter 23, pages 387–394. ISBN 0-8247-6007-7 (hardcover). LCCN TS156.8 .S7537 2000.

Longford:1999:SVS

[LN99c] Nicholas T. Longford and John A. Nelder. Statistics versus statistical science in the regulatory process. Statistics in Medicine, 18 (17–18):2311–2320, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2000:HAC

[LN00a] Y. Lee and J. A. Nelder. HGLMs for analysis of correlated nonnormal data. In Bethlehem and van der Heijden [BvdH00], pages 97–107. ISBN 3-7908-1326-5 (paperback), 3-642-57678-8 (e-book). LCCN QA276.4 .C667 2000.

Lee:2000:RBD

[LN00b] Y. Lee and J. A. Nelder. The relationship between doubleexponential families and extended quasi-likelihood families, with application to modelling Geissler’s human sex ratio data. Applied Statistics, 49(3):413–419, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2000:TWM

[LN00c] Y. Lee and J. A. Nelder. Two ways of modelling overdispersion in non-normal data. Applied Statistics, 49(4):591–598, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2001:HGL

[LN01] Youngjo Lee and John A. Nelder. Hierarchical generalised linear models: a synthesis of generalised linear models, random-effect models and structured dispersions. Biometrika, 88(4):987–1006, 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Lee:2002:AUD

[LN02] Youngjo Lee and John A. Nelder. Analysis of ulcer data using hierarchical generalized linear models. Statistics in Medicine, 21 (2):191–202, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2003:ERE

[LN03a] Youngjo Lee and John A. Nelder. Extended-REML estimators. Journal of Applied Statistics, 30(8):845–856, 2003. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2003:FPD

[LN03b] Youngjo Lee and John A. Nelder. False parsimony and its detection with GLMs. Journal of Applied Statistics, 30(5):477–483, 2003. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2004:CMM

[LN04] Youngjo Lee and John A. Nelder. Conditional and marginal models: another view. Statistical Science, 19(2):219–238, 2004. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Lee:2005:LRE

[LN05] Youngjo Lee and John A. Nelder. Likelihood for random-effect models. SORT, 29(2):141–182, 2005. With a discussion and a rejoinder by the authors.

Lee:2006:DHG

[LN06a] Youngjo Lee and John A. Nelder. Double hierarchical generalized linear models. Applied Statistics, 55(2):139–185, 2006. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2006:FAR

[LN06b] Youngjo Lee and John A. Nelder. Fitting via alternative randomeffect models. Statistics and Computing, 16(1):69–75, 2006. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Lee:2009:LIM

[LN09a] Youngjo Lee and John A. Nelder. Likelihood inference for models with unobservables: another view. Statistical Science, 24(3):255– 269, 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Lee:2009:RLI

[LN09b] Youngjo Lee and John A. Nelder. Rejoinder: Likelihood inference for models with unobservables another view [MR2757427; MR2757428; MR2757430; MR2757429]. Statistical Science, 24(3): 294–302, 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Lee:2007:LPS

[LNN07] Youngjo Lee, John A. Nelder, and Maengseok Noh. H-likelihood: problems and solutions. Statistics and Computing, 17(1):49–55, 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Lee:2006:GLM

[LNP06] Youngjo Lee, John A. Nelder, and Yudi Pawitan. Generalized Linear Models with Random Effects, volume 106 of Monographs on Statistics and Applied Probability. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-631-5; 1-58488-631-5. x + 396 pp. Unified analysis via H-likelihood, With 1 CD-ROM (Windows).

Lee:2011:HQI

[LNP11] Youngjo Lee, John A. Nelder, and Heejin Park. HGLMs for quality improvement. Applied Stochastic Models in Business and Industry, 27(3):315–328, 2011.

Lee:2017:GLM

[LNP17] Youngjo Lee, John A. Nelder, and Yudi Pawitan. Generalized Linear Models with Random Effects, volume 153 of Monographs on Statistics and Applied Probability. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2017. ISBN 1-4987-2061-7. xix + 446 pp. Unified analysis via H-likelihood.

Locsi:2013:HVN

[Lóc13] Levente Lócsi. A hyperbolic variant of the Nelder–Mead simplex method in low dimensions. Acta Universitatis Sapientiae. Mathematica. An International Scientific Journal of Sapientia University, 5(2):169–183, 2013.

Lagarias:2012:CRN

[LPW12] Jeffrey C. Lagarias, Bjorn Poonen, and Margaret H. Wright. Convergence of the restricted Nelder–Mead algorithm in two dimensions. SIAM Journal on Optimization, 22(2):501–532, ???? 2012. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Liang:1997:DFD

[LR97] Kung-Yee Liang and Paul J. Rathouz. Discussion on A. F. Desmond’s paper: Optimal estimating functions, quasilikelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):113–116, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900148. See [Des97a].

Lagarias:1998:CPN

[LRWW98] Jeffrey C. Lagarias, James A. Reeds, Margaret H. Wright, and Paul E. Wright. Convergence properties of the Nelder–Mead simplex method in low dimensions. SIAM Journal on Optimization, 9(1):112–147, October/December 1998. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/30347.

Lagarias:1999:CPN

[LRWW99] Jeffrey C. Lagarias, James A. Reeds, Margaret H. Wright, and Paul E. Wright. Convergence properties of the Nelder–Mead simplex method in low dimensions. SIAM Journal on Optimization, 9(1):112–147, 1999. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Li:1995:OFL

[LTJ95] Wei Li, Qunhua Tan, and Hartmut Janocha. Optimizing fuzzy logic controller using Nelder and Mead’s simplex algorithm. Control Theory & Applications. Kongzhi Lilun yu Yingyong, 12(5): 616–622, 1995. ISSN 1000-8152.

Liu:2012:NHN

[LY12] An Liu and Ming-Ta Yang. A new hybrid Nelder–Mead particle swarm optimization for coordination optimization of directional overcurrent relays. Mathematical Problems in Engineering, pages 1–18, 2012.

Mastorakis:2006:UOD

[Mas06] Nikos E. Mastorakis. Unstable ordinary differential equations: solution via genetic algorithms and the method of Nelder–Mead. WSEAS Transactions on Mathematics, 5(12):1276–1281, 2006. ISSN 1109-2769 (print), 2224-2880 (electronic).

Mastorakis:2007:SVD

[Mas07] Nikos E. Mastorakis. The singular value decomposition (SVD) in tensors (multidimensional arrays) as an optimization problem. Solution via genetic algorithms and method of Nelder–Mead. WSEAS Transactions on Systems, 6(1):17–23, 2007. ISSN 1109-2777 (print), 2224-2678 (electronic).

Mastorakis:2009:GAN

[Mas09a] Nikos E. Mastorakis. Genetic algorithms with Nelder–Mead optimization in the variational methods of boundary value problems. WSEAS Transactions on Mathematics, 8(3):107–116, 2009. ISSN 1109-2769 (print), 2224-2880 (electronic).

Mastorakis:2009:SSM

[Mas09b] Nikos E. Mastorakis. Solution of the Schrödinger–Maxwell equations via finite elements and genetic algorithms with Nelder–Mead. WSEAS Transactions on Mathematics, 8(4):169–176, 2009. ISSN 1109-2769 (print), 2224-2880 (electronic).

McKinnon:1998:CNM

[McK98] K. I. M. McKinnon. Convergence of the Nelder–Mead simplex method to a nonstationary point. SIAM Journal on Optimization, 9(1):148–158, October/December 1998. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/30348.

McKinnon:1999:CNM

[McK99] K. I. M. McKinnon. Convergence of the Nelder–Mead simplex method to a nonstationary point. SIAM Journal on Optimization, 9(1):148–158, 1999. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Mehta:2020:INM

[Meh20] V. K. Mehta. Improved Nelder–Mead algorithm in high dimensions with adaptive parameters based on Chebyshev spacing points. Engineering Optimization, 52(10):1814–1828, 2020. CO-DEN EGOPAX. ISSN 0305-215X.

McCullagh:2011:JAN

[MGP11] Peter McCullagh, Bob Gilchrist, and Roger Payne. John Ashworth Nelder, 1924–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):499–502, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

http://www.jstor.org/stable/23014410.

Moraglio:2010:GGN

[MJ10] Alberto Moraglio and Colin G. Johnson. Geometric generalization of the Nelder–Mead algorithm. In Evolutionary computation in combinatorial optimization, volume 6022 of Lecture Notes in Comput. Sci., pages 190–201. Springer, Berlin, 2010. ISBN 3-642-12138-1; 3-642-12138-1.

Majhi:2026:PGA

[MM26] Neha Majhi and Rajashree Mishra. Partitioned genetic algorithm with Nelder–Mead for real-coded problem-solving and it’s application. In Proceedings of the International Conference on Computational Mathematics and Applications, volume 1520 of Lect. Notes Netw. Syst., pages 169–182. Springer, Singapore, 2026. ISBN 981-9689-86-4; 981-9689-87-2.

Milton:1969:SC

[MN69] Roy C. Milton and John Ashworth Nelder, editors. Statistical Computation. Academic Press, New York, NY, USA, 1969. ISBN 0-12-498150-X (hardcover), 1-4832-5802-5 (e-book). xii + 462 pp.

LCCN QA276.4 C74 1969. URL https://www.sciencedirect.
com/book/edited-volume/9780124981508/statistical-computation.

McCullagh:1983:GLM

[MN83] P. McCullagh and J. A. Nelder. Generalized Linear Models. Monographs on Statistics and Applied Probability. Chapman and Hall, Ltd., London, UK, 1983. ISBN 0-412-23850-0. xiii + 261 pp. LCCN H61.25; QA276.M38.

McCullagh:1989:GLM

[MN89] P. McCullagh and J. A. Nelder. Generalized Linear Models. Monographs on Statistics and Applied Probability. Chapman and Hall, Ltd., London, UK, second edition, 1989. ISBN 0-412-31760-5. xix + 511 pp.

Milton:2014:SC

[MN14] Roy C. Milton and John A. Nelder, editors. Statistical Computation. Elsevier Science and Technology, 2014. ISBN 1-4832-5802-5. LCCN QA276.4 S738 1969.

Nelder:1960:ASY

[NABS60] J. A. Nelder, R. B. Austin, J. K. A. Bleasdale, and P. J. Salter. An approach to the study of yearly and other variation in crop yields. Journal of Horticultural Science, 35(1):73–82, 1960. CODEN JH-

SOFF. ISSN 2582-4899. URL https://www.tandfonline.com/
doi/abs/10.1080/00221589.1960.11513972.

Nair:1992:TPD

[NAM^{+}92] Vijayan N. Nair, Bovas Abraham, Jock MacKay, John A. Nelder, George Box, Madhav S. Phadke, Raghu N. Kacker, Jerome Sacks, William J. Welch, Thomas J. Lorenzen, Anne C. Shoemaker, Kwok L. Tsui, James M. Lucas, Shin Taguchi, Raymond H. Myers, G. Geoffrey Vining, and C. F. Jeff Wu. Taguchi’s parameter design: A panel discussion. Technometrics, 34(2):127–161, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269231. Edited

by Vijayan N. Nair.

Nelder:1982:GLMa

[NB82a] J. A. Nelder and R. J. Baker. Generalized linear models. In Ferber et al. [FKJ^{+}89], pages 343–348. ISBN 0-471-05546-8 (vol. 1), 0-471-05547-6 (vol. 2), 0-471-05549-2 (vol. 3), 0-471-05551-4 (vol. 4), 0-471-05552-2 (vol. 5), 0-471-05553-0 (vol. 6), 0-471-05555-7 (vol. 7), 0-471-05556-5 (vol. 8), 0-471-81274-3 (vol. 10), 0-471-85474-3 (vol. 9). LCCN QA276.14 .E53 1981.

Nelder:1982:G

[NB82b] J. A. Nelder and R. J. Baker. GLIM. In Ferber et al. [FKJ^{+}89], pages 439–442. ISBN 0-471-05546-8 (vol. 1), 0-471-05547-6 (vol. 2), 0-471-05549-2 (vol. 3), 0-471-05551-4 (vol. 4), 0-471-05552-2 (vol. 5), 0-471-05553-0 (vol. 6), 0-471-05555-7 (vol. 7), 0-471-05556-5 (vol. 8), 0-471-81274-3 (vol. 10), 0-471-85474-3 (vol. 9). LCCN QA276.14 .E53 1981.

Nelder:1985:SSP

[NB85] J. A. Nelder and R. J. Baker. Statistical software; progress and prospects. In Billard [Bil85], pages 33–37. ISBN 0-444-87725-8 (paperback). LCCN QA276.4 .S95 1984.

Nelder:1966:MRD

[NBCS66] J. A. Nelder, G. Berry, T. J. Cleaver, and P. J. Salter. Methods of recording data in the laboratory and field. Exp. Agr., 2(??):69–80, ???? 1966.

Nelder:1967:SPE

[NC67] J. A. Nelder and B. E. Cooper. Statistical programming: Epilogue. Applied Statistics, 16(2):149–151, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1971:SCI

[NC71] J. A. Nelder and B. E. Cooper. Statistical computing 1970: Input/ output in statistical programming. Applied Statistics, 20(1):56– 73, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1970:CDA

[NCC^{+}70] J. A. Nelder, B. E. Cooper, J. M. Craddock, J. C. Gower, P. J. Harrison, and I. D. Hill. The construction and description of algorithms. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):255–260, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.jstor.

org/stable/2986837.

Nelder:1951:NSI

[Nel51] J. A. Nelder. A note on the statistical independence of quadratic forms in the analysis of variance. Biometrika, 38:482–483, 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Nelder:1952:SGF

[Nel52] J. A. Nelder. Some genotypic frequencies and variance components occurring in biometrical genetics. Heredity, 6(??):387– 394, December 1, 1952. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://www.nature.com/

articles/hdy195247.

Nelder:1953:SMB

[Nel53] J. A. Nelder. Statistical models in biometrical genetics. Heredity, 7(??):111–119, April 1, 1953. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://www.nature.com/

articles/hdy195311.

Nelder:1954:INC

[Nel54a] J. A. Nelder. The interpretation of negative components of variance. Biometrika, 41:544–548, 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Nelder:1954:NMP

[Nel54b] J. A. Nelder. A note on missing plot values. Biometrics, 10(3): 400–401, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/3001594.

Nelder:1956:BRB

[Nel56a] J. A. Nelder. Book review: Experimental Design, Theory and Application, by Walter T. Federer. Journal of the Royal Statistical Society. Series A (General), 119(3):340–341, ???? 1956. CO-DEN JSSAEF. ISSN 0035-9238. URL https://www.jstor.org/

stable/2342744.

Nelder:1956:NPE

[Nel56b] J. A. Nelder. Notes on the planning and executing of field experiments. N. A. A. S. Quarterly Review, 34(??):139–146, ???? 1956. CODEN NAQRA4. ISSN 0374-9487.

Nelder:1959:NVR

[Nel59] J. A. Nelder. National Vegetable Research Station new laboratory building. Nature, 184(4696):1368–1369, October 31, 1959. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL https://www.nature.com/articles/1841368a0.

Nelder:1960:EVC

[Nel60] J. A. Nelder. The estimation of variance components in certain types of experiment on quantitative genetics. In O. Kempthorne, editor, Biometrical Genetics, pages 139–158. Pergamon, New York, NY, USA, 1960.

Nelder:1961:FGL

[Nel61a] J. A. Nelder. The fitting of a generalization of the logistic curve. Biometrics, 17:89–110, 1961. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

Nelder:1961:NSG

[Nel61b] J. A. Nelder. A note on some growth patterns in a simple theoretical organism. Biometrics, 17:220–228, 1961. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

Nelder:1962:NAF

[Nel62a] J. A. Nelder. 182. Note: an alternative form of a generalized logistic equation. Biometrics, 18(4):614–616, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527907.

Nelder:1962:NKS

[Nel62b] J. A. Nelder. New kinds of systematic designs for spacing experiments. Biometrics, 18(3):283–307, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527473.

Nelder:1962:QGG

[Nel62c] J. A. Nelder. Quantitative genetics and growth analysis. Statistical genetics and plant breeding. Washington National Academy of Sciences, ??(??):445–545, ???? 1962.

Nelder:1962:SEH

[Nel62d] J. A. Nelder. A statistical examination of the Hastings Rarities. British Birds, 55(??):283–297, August 1962. URL https:

//www.britishbirds.co.uk/journal/article/statistical-
examination-hastings-rarities.

Nelder:1963:BRB

[Nel63a] J. A. Nelder. Book review: The Essence of Biometry, by John Stanley. Journal of the Royal Statistical Society. Series A (General), 126(4):600–601, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.org/stable/2982590.

Nelder:1963:ICF

[Nel63b] J. A. Nelder. Identification of contrasts in fractional replicates of 2^{n} experiments. Applied Statistics, 12(1):38–43, March 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1963:URS

[Nel63c] J. A. Nelder. The use of response surfaces in the interpretation of groups of experiments. In ????, editor, 5th International Biometric Conference, Cambridge, September 10–14, 1963: preprints, page ?? ????, ????, 1963. LCCN QH323.5 .I56 1963.

Nelder:1963:YDR

[Nel63d] J. A. Nelder. Yield-density relations and Jarvis’s lucerne data. Journal of Agricultural Science, 61(3):427–429, ???? 1963. CODEN JASIAB. ISSN 0021-8596 (print), 1469-5146 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-agricultural-science/article/yielddensity-
relations-and-jarviss-lucerne-data/23F06B5DD52CAD6E8E88A7EDA5D8CAB9.

Nelder:1964:BRE

[Nel64a] J. A. Nelder. Book review: Les Essais thérapeutiques cliniques, by D. Schwartz, R. Flamant, J. Lellouch, C. Rouquette. Biometrics, 20(2):376, June 1964. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/

2528410. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Nelder:1964:EN

[Nel64b] J. A. Nelder. Editorial note. Biometrics, 20(4):883, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528138.

Nelder:1965:AREa

[Nel65a] J. A. Nelder. The analysis of randomized experiments with orthogonal block structure. I. Block structure and the null analysis of variance. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 283:147–162, 1965. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Nelder:1965:AREb

[Nel65b] J. A. Nelder. The analysis of randomized experiments with orthogonal block structure. II. Treatment structure and the general analysis of variance. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 283:163–178, 1965. CO-DEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Nelder:1966:EEH

[Nel66a] J. A. Nelder. Evolutionary experimentation in horticulture. Journal of the Science of Food and Agriculture, 17(1):7–9, January 1966. CODEN JSFAAE. ISSN 0022-5142 (print), 1097-

0010 (electronic). URL https://scijournals.onlinelibrary.
wiley.com/doi/10.1002/jsfa.2740170102.

Nelder:1966:IPU

[Nel66b] J. A. Nelder. Inverse polynomials, a useful group of multi-factor response functions. Biometrics, 22(1):128–141, March 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528220.

Nelder:1967:ACA

[Nel67] J. A. Nelder. The application of computers to agricultural and horticultural research. Agric. Prog., 42(??):7–23, ???? 1967. Middleton Memorial Lecture.

Nelder:1968:AAM

[Nel68a] J. A. Nelder. Algorithms: Algorithm AS 8: Main effects from a multi-way table. Applied Statistics, 17(3):277–279, September 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.edu/apstat/8.

Nelder:1968:AAC

[Nel68b] J. A. Nelder. Algorithms: Algorithm AS 9: Construction of additive table. Applied Statistics, 17(3):279–283, September 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

URL http://lib.stat.cmu.edu/apstat/9.

Nelder:1968:BR

[Nel68c] J. A. Nelder. Book review. The Computer Journal, 10(4):367, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/367.full.pdf+html.

Nelder:1968:CIG

[Nel68d] J. A. Nelder. The combination of information in generally balanced designs. Journal of the Royal Statistical Society. Series B (Methodological), 30:303–311, 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://links.jstor.org/sici?sici=0035-9246(
1968)30:2<303:TCOIIG>2.0.CO;2-D&origin=MSN.

Nelder:1968:RMB

[Nel68e] J. A. Nelder. Regression, model-building and invariance. Journal of the Royal Statistical Society. Series A (General), 131(3):303– 329, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL https:

//www.jstor.org/stable/2343525.

Nelder:1968:WRQ

[Nel68f] J. A. Nelder. Weighted regression, quantal response data, and inverse polynomials. Biometrics, 24(4):979–985, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528884.

Nelder:1969:DDS

[Nel69a] J. A. Nelder. The description of data structures for statistical computing. In Milton and Nelder [MN69], pages 13–36. ISBN 0-12-498150-X (hardcover), 1-4832-5802-5 (e-book). LCCN QA276.4

C74 1969.  URL https://www.sciencedirect.com/book/
edited-volume/9780124981508/statistical-computation.

Nelder:1969:SAA

[Nel69b] J. A. Nelder. Statistical algorithms: Algorithm AS 20: The efficient formation of a triangular array with restricted storage for data. Applied Statistics, 18(2):203–206, June 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://lib.stat.cmu.edu/apstat/20.

Nelder:1971:BRB

[Nel71a] J. A. Nelder. Book review: Statistics In Biology, by Chester I. Bliss. Journal of the Royal Statistical Society. Series A (General), 134(2):246–247, ???? 1971. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/2343879.

Nelder:1971:SCS

[Nel71b] J. A. Nelder. Statistical computing 1970: Statistical computing and computer languages. Applied Statistics, 20(1):25–32, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1971:TVC

[Nel71c] J. A. Nelder. [Topics in variance component estimation]. Biometrics, 27(3):750, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528616.

Nelder:1972:STR

[Nel72a] J. Nelder. Statistical training and research/formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 40(3):384–388, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402476.

Nelder:1972:BRBa

[Nel72b] J. A. Nelder. Book review: Analysis and Design of Certain Quantitative Multiresponse Experiments by S. N. Roy; R. Gnanadesikan; J. N. Srivastava. International Statistical Review = Revue Internationale de Statistique, 40(2):256, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402776.

Nelder:1972:BRBb

[Nel72c] J. A. Nelder. Book review: Probability, Statistics and Data Analysis, by Oscar Kempthorne; Leroy Folks. Journal of the Royal Statistical Society. Series A (General), 135(3):430–431, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2344621.

Nelder:1972:MB

[Nel72d] J. A. Nelder. Mathematics for biologists. Bulletin of the Institute of Mathematics and its Applications, 8(??):217–219, ???? 1972. CODEN IMTABW. ISSN 0950-5628.

Nelder:1972:SAS

[Nel72e] J. A. Nelder. Summary and assessment: a statistician’s point of view. In Jeffers [Jef72], pages 367–373. ISBN 0-632-08740-4. LCCN QH541.15M3 1972.

Nelder:1974:GSS

[Nel74a] J. A. Nelder. GenStat: a statistical system. In Bruckmann et al. [BFS74], page ?? ISBN 3-7908-0148-8. ISSN 0253-018X. LCCN QA276.4 .C18 1974.

Nelder:1974:LLM

[Nel74b] J. A. Nelder. Log linear models for contingency tables: a generalization of classical least squares. Applied Statistics, 23:323–329, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1974:SPI

[Nel74c] J. A. Nelder. Statistical packages. Introductory remarks. Bulletin of the Institute of Mathematics and its Applications, 10(5–6):165– 166, ???? 1974. CODEN IMTABW. ISSN 0950-5628.

Nelder:1974:UGE

[Nel74d] J. A. Nelder. A user’s guide to the evaluation of statistical packages and systems. International Statistical Review = Revue Internationale de Statistique, 42(3):291–298, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402986.

Nelder:1975:AWP

[Nel75a] J. A. Nelder. Announcement by the working party on statistical computing: GLIM (generalized linear interactive modelling program). Applied Statistics, 24(2):259–261, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1975:BRB

[Nel75b] J. A. Nelder. Book review: The Analysis of Categorical Data, by R. L. Plackett. Journal of the Royal Statistical Society. Series A (General), 138(1):99–100, ???? 1975. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.org/stable/2345253.

Nelder:1975:IG

[Nel75c] J. A. Nelder. An introduction to GenStat. The Mathematical Scientist, 1(1):57–58, ???? 1975. ISSN 0312-3685 (print), 1475-6080 (electronic).

Nelder:1975:RWM

[Nel75d] J. A. Nelder. Robert William MacLagan Wedderburn, 1947–1975. Journal of the Royal Statistical Society. Series A (General), 138 (4):587, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2345239.

Nelder:1975:STR

[Nel75e] J. A. Nelder. Statistical training and research/formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 43(2):227–232, August 1975. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402904.

Nelder:1975:CB

[Nel75f] John A. Nelder. Computers in Biology. The Wykeham science series. Wykeham Publications, London, UK, 1975. ISBN 0-387-91120-0. viii + 158 pp. LCCN QH324.2 .N44 1975.

Nelder:1976:IPN

[Nel76a] J. A. Nelder. Intelligent programmes, the next stage in statistical computing. In Barra et al. [BS^{+}97], page ?? ISBN 0-7204-0751-6. LCCN QA276.A1 E89 1976.

Nelder:1976:SAA

[Nel76b] J. A. Nelder. Statistical algorithms: Algorithm AS 96: a simple algorithm for scaling graphs. Applied Statistics, 25(1):94–96, March 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-

9876 (electronic). URL http://lib.stat.cmu.edu/apstat/96.

Nelder:1977:BRB

[Nel77a] J. A. Nelder. Book review: Theory and Application of the Linear Model, by Franklin A. Graybill. Journal of the Royal Statistical Society. Series A (General), 140(3):384–385, ???? 1977. CODEN

JSSAEF. ISSN 0035-9238. URL http://www.jstor.org/stable/
2344939.

Nelder:1977:MDC

[Nel77b] J. A. Nelder. Multi-dimensional contingency table with one factor as a response. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):41–42, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988218.

Nelder:1977:MKS

[Nel77c] J. A. Nelder. My kind of statistics. Bulletin in Applied Statistics (BIAS), 4(2):77–81, 1977. CODEN ???? ISSN ????

Nelder:1977:RLM

[Nel77d] J. A. Nelder. A reformulation of linear models. Journal of the Royal Statistical Society. Series A (General), 140(1):48–76, 1977. CODEN JSSAEF. ISSN 0035-9238. With discussion.

Nelder:1978:BRB

[Nel78a] J. A. Nelder. Book review: Computational Methods for Data Analysis, by John M. Chambers. Journal of the Royal Statistical Society. Series A (General), 141(4):550–551, ???? 1978. CODEN JS-

SAEF. ISSN 0035-9238. URL http://www.jstor.org/stable/
2344494.

Nelder:1978:FSS

[Nel78b] J. A. Nelder. The future of statistical software. In Corsten and Hermans [CH78], pages 11–19. ISBN 3-7908-0196-8 (paperback). ISSN 0253-018X. LCCN QA276.4 .C192.

Nelder:1978:RMB

[Nel78c] J. A. Nelder. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73 (361):89–90, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286523.

Nelder:1979:BRR

[Nel79a] J. A. Nelder. Book review: Regression Analysis by Example, by S. Chatterjee, B. Price. Biometrics, 35(1):355–356, March 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529957.

Nelder:1979:BRS

[Nel79b] J. A. Nelder. Book review: Statistical Analysis, a Computer Oriented Approach, by A. A. Afifi, S. P. Azen. Biometrics, 35(3):703, September 1979. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/
2530266.

Nelder:1979:BRB

[Nel79c] J. A. Nelder. Book review: The General Linear Model, Data Analysis in the Social and Behavioural Sciences, by R. L. Horton. Journal of the Royal Statistical Society. Series A (General), 142(3):385, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2982495.

Nelder:1979:BRA

[Nel79d] J. A. Nelder. Book reviews: The Analysis of Cross-Tabulated Data, by Graham J. G. Upton. Applied Statistics, 28(3):306–307, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1979:EDS

[Nel79e] J. A. Nelder. Experimental design and statistical evaluation. In Fosdick [Fos79], pages 309–315. ISBN 0-444-85330-8 (hardcover). LCCN QA297 .I18 1978.

Nelder:1980:BRS

[Nel80a] J. A. Nelder. Book review: Statistical Computing, by W. J. Kennedy, Jr., J. E. Gentle. Biometrics, 36(4):743–744, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556137.

Nelder:1980:BRB

[Nel80b] J. A. Nelder. Book review: Statistical Methods for Digital Computers by K. Enslein; A. Ralston; H. S. Wilf. International Statistical Review = Revue Internationale de Statistique, 48(1):120–121, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823

(electronic). URL http://www.jstor.org/stable/1402419.

Nelder:1980:IWL

[Nel80c] J. A. Nelder. Iterative weighted least squares; an algorithm for many occasions. In Diday et al. [DLPT80], pages 75–81. ISBN 0-444-86005-3 (hardcover). LCCN QA278 .I56 1979.

Nelder:1980:QSS

[Nel80d] J. A. Nelder. Les qualités souhaitables dans des systèmes statistiques en se servant de GenStat comme référence. (French) [Desirable qualities in statistical systems using GenStat as a reference]. Statistique et Analyse des Données, 5(3):17–27, ???? 1980. ISSN 0750-7364 (print), 2400-4766 (electronic). URL https://

www.numdam.org/item/SAD_1980__5_3_17_0.pdf.

Nelder:1982:AGL

[Nel82a] J. A. Nelder. Analysis of generalized linear models using GLIM. Metron, 40(??):243–248, ???? 1982. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic).

Nelder:1982:FDR

[Nel82b] J. A. Nelder. Factorial designs: [review of book by B. L. Raktoe, A. Hedayat and W. T. Federer, Wiley, New York, 1981, 209 pages. European Journal of Operational Research, 10(1):119–120, May 1982. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Nelder:1982:GLMb

[Nel82c] J. A. Nelder. Generalized Linear Models; a useful synthesis in statistics. I.H.S. Journal, 5(??):191–201, ???? 1982.

Nelder:1982:LMN

[Nel82d] J. A. Nelder. Linear models and nonorthogonal data. Utilitas Mathematica, 21:141–152, 1982. CODEN UTMADA. ISSN 0315-3681. Special issue dedicated to Frank Yates on the occasion of his eightieth birthday, Vol. B.

Nelder:1983:RMO

[Nel83a] J. A. Nelder. The role of models in official statistics. In Eurostat Seminar on recent developments in the analysis of large scale data sets, 16–18 November 1982, page ???? Secretariate, P B 1907-Luxembourg, 1983. Item No. 01.

Nelder:1983:PPC

[Nel83b] John A. Nelder. Presentation of papers at conferences. Biometrics, 39(3):799, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531116.

Nelder:1984:DSN

[Nel84a] J. A. Nelder. Do statisticians need special interactive languages? In Diday [Did84], pages 511–516. ISBN 0-444-87555-7 (hardcover). LCCN QA278 .I56 1981.

Nelder:1984:PPP

[Nel84b] J. A. Nelder. Present position and potential developments: Some personal views: Statistical computing. Journal of the Royal Statistical Society. Series A (General), 147(2):151–160, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2981672.

Nelder:1985:QLG

[Nel85a] J. A. Nelder. Quasi-likelihood and GLIM. In Generalized linear models (Lancaster, 1985), volume 32 of Lect. Notes Stat., pages 120–127. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-96224-7.

Nelder:1985:SMQ

[Nel85b] J. A. Nelder. Statistical models for qualitative data. In Nijkamp et al. [NLW85], pages 31–38. ISBN 90-247-3124-0. LCCN QA278.2

.N38 1983. URL http://www.gbv.de/dms/tib-ub-hannover/
01664381X.pdf.

Nelder:1985:AIS

[Nel85c] John A. Nelder. An alternative interpretation of the singularvalue decomposition in regression. The American Statistician, 39 (1):63–64, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683911.

Nelder:1986:GAM

[Nel86a] J. A. Nelder. [Generalized additive models]: Comment. Statistical Science, 1(3):312, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013606.

Nelder:1986:SST

[Nel86b] J. A. Nelder. Statistics, science and technology. Journal of the Royal Statistical Society. Series A (General), 149(2):109–121, ???? 1986. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981525.

Nelder:1987:AIG

[Nel87a] J. A. Nelder. Artificial intelligence and generalized linear modelling: An expert system for GLIM. In Phelps [Phe87], pages 36–44. ISBN 0-291-39743-3 (paperback). LCCN Q334 .I548 1987.

Nelder:1987:BRS

[Nel87b] J. A. Nelder. Book review: Statistics in Ornithology, by B. J. T. Morgan, P. M. North. Biometrics, 43(1):253–254, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531971.

Nelder:1987:DWA

[Nel87c] J. A. Nelder. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):930–931, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350480.

Nelder:1988:HSS

[Nel88a] J. A. Nelder. How should the statistical expert system and its user see each other? In COMPSTAT 1988 (Copenhagen, 1988), pages 107–116. Physica, Heidelberg, West Germany, 1988. ISBN 3-7908-0411-8.

Nelder:1988:RES

[Nel88b] J. A. Nelder. The role of expert systems in statistics. In Faulbaum and Uehlinger [FU88], pages 175–182. ISBN 3-437-50320-0. LCCN QA276.4 .K66 1987.

Nelder:1992:CPS

[Nel88c] J. A. Nelder. Comment on paper by Streitberg, B. Statistical Software Newsletter. Computational Statistics & Data Analysis, 14(??):73–??, ???? 1992 (or 1988??). CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Nelder:1989:DPP

[Nel89] J. A. Nelder. Discussion of the paper by Professors Godambe and Thompson. Journal of Statistical Planning and Inference, 22(2): 153–172, June 1989. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375889901079. See [GT89].

Nelder:1990:NPL

[Nel90a] J. A. Nelder. Nearly parallel lines in residual plots. The American Statistician, 44(3):221–222, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685341.

Nelder:1990:ODB

[Nel90b] J. A. Nelder. Obituary: Dr. Brian Edward Cooper, 1935–90. Journal of the Royal Statistical Society. Series A (General), 153(3):397, 1990. CODEN JSSAEF. ISSN 0035-9238.

Nelder:1991:GKB

[Nel91] John A. Nelder. GLIMPSE, a knowledge-based front end for GLIM. In Buja and Tukey [BT91], pages 125–131. ISBN 0-387-97633-7. LCCN QA276.4 .C59 1991.

Nelder:1992:CSA

[Nel92a] J. A. Nelder. The computer as statistical assistant. In Gaffney and Houstis [GH92], pages 69–75. ISBN 0-444-89176-5 (paperback). ISSN 0926-5473. LCCN QA76.6 .I1782 1991. URL http://www.

gbv.de/dms/bowker/toc/9780444891761.pdf;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0746.00068.

Nelder:1992:JMM

[Nel92b] J. A. Nelder. Joint modelling of mean and dispersion. In Statistical Modelling, pages 263–272. Elsevier Science Publishers B.V., ????, 1992.

Nelder:1992:MS

[Nel92c] J. A. Nelder. Modelling in statistics. Mathematical and Computer Modelling, 16(4):131–136, April 1992. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
089571779290040R.

Nelder:1992:MCT

[Nel92d] J. A. Nelder. Models for curves through the origin. Journal of Applied Statistics, 19(4):551–552, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nelder:1992:SPU

[Nel92e] John A. Nelder. Statistical packages and unbalanced data. Computational Statistics & Data Analysis, 14(3):403–406, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016794739290050P.

Nelder:1993:ASC

[Nel93a] J. A. Nelder. Aspects of statistical computing, past, present and future. Statistica Neerlandica, 47(1):3–8, March 1993. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic).

URL https://onlinelibrary.wiley.com/doi/epdf/10.1111/
j.1467-9574.1993.tb01402.x.

Nelder:1993:MIA

[Nel93b] J. A. Nelder. The most important areas of statistical research in the next ten years. Statistics and Computing, 3(4): 202–203, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00141783.

Nelder:1994:AVS

[Nel94a] J. A. Nelder. An alternative view of the splicing data. Applied Statistics, 43(3):469–476, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1994:BRB

[Nel94b] J. A. Nelder. Book review: Random Coefficient Models, by N. T. Longford. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):507, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983541.

Nelder:1994:RAP

[Nel94c] J. A. Nelder. A re-analysis of the pump-failure data. Comment on: “Conjugate likelihood distributions” [Scand. J. Statist. 20 (1993), no. 2, 147–156; MR1229290 (94g:62007)] by E. I. George, U. E. Makov and A. F. M. Smith. Scandinavian Journal of Statistics. Theory and Applications, 21(2):187–191, 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). With a reply by George, Makov and Smith.

Nelder:1994:SLM

[Nel94d] J. A. Nelder. The statistics of linear models: back to basics. Statistics and Computing, 4(4):221–234, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF00156745. See comments [Nel95c, Ait95, Gow95, Van95, RTW95, Sea95] and rejoinder [Nel95b].

Nelder:1995:GLM

[Nel95a] J. A. Nelder. Generalized linear models: a powerful tool for the analysis of quality-improvement experiments. In ????, editor, Proceedings of the International Conferences on Statistical Methods and Statistical Computing for Quality and Productivity Improvement, Seoul, page ?? ????, ????, 1995.

Nelder:1995:RCS

[Nel95b] J. A. Nelder. Rejoinder to comments on ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):109–111, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-

1375 (electronic). URL http://link.springer.com/article/
10.1007/BF00143940. See [Nel94d].

Nelder:1995:SLM

[Nel95c] J. A. Nelder. The statistics of linear models: back to basics. Statistics and Computing, 5(2):i, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00143933. See

[Nel94d].

Nelder:1997:FY

[Nel97a] J. A. Nelder. Frank Yates. In Johnson and Kotz [JK97], pages 347– 349. ISBN 0-471-16381-3 (paperback), 1-118-15071-6 (e-book), 1-118-15072-4, 1-283-28289-5. LCCN QA276.156 .L43 1997.

Nelder:1997:GMM

[Nel97b] J. A. Nelder. The great mixed-model muddle is alive and flourishing — alas! Food Qual. Prefer., 9(??):157–159, ???? 1997.

Nelder:1997:CDF

[Nel97c] John Nelder. Contribution to discussion on A. F. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1): 108, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900124. See [Les97].

Nelder:1998:SSS

[Nel98a] J. A. Nelder. From statistics to statistical science. Journal of the Royal Statistical Society. Series D (The Statistician), 48(??):269– ??, ???? 1998. ISSN 0039-0526 (print), 1467-9884 (electronic). Not in jrss-d-1990.bib and article containing page 269 does not have Nelder comments.

Nelder:1998:LEB

[Nel98b] J. A. Nelder. Letter to the Editor: A Bayesian analysis of regression models with continuous errors with application to longitudinal studies, by L. D. Broemeling and P. Cook, Statistics in Medicine, 16, 321–332 (1997). Statistics in Medicine, 17(2):241, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [BC97].

Nelder:1998:LEI

[Nel98c] J. A. Nelder. Letters to the Editors: The importance of marginality rules. Applied Statistics, 47(3):447–448, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

URL http://www.blackwellpublishers.co.uk/asp/journal.
asp?ref=0035-9254&src=ard&aid=122&iid=3&vid=47.

Nelder:1998:LCM

[Nel98d] John A. Nelder. A large class of models derived from generalized linear models. Statistics in Medicine, 17(23):2747–2753, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelder:1998:STR

[Nel98e] John A. Nelder. The selection of terms in response-surface models: How strong is the weak-heredity principle? The American Statistician, 52(4):315–318, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/nelder.html;   http://www.jstor.org/stable/2685433.

Nelder:1999:SSS

[Nel99] John A. Nelder. From statistics to statistical science. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2): 257–269, ???? 1999. CODEN ???? ISSN 0039-0526 (print),

1467-9884 (electronic). URL http://www.jstor.org/stable/
2681191.

Nelder:2000:FMR

[Nel00a] J. A. Nelder. Functional marginality and response-surface fitting. Journal of Applied Statistics, 27(1):109–112, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (elec-

tronic). URL http://www.catchword.co.uk/cgi-bin/cgi?ini=
carfax&body=linker&reqidx=/catchword/carfax/13600532/
v27n1/s8/p109.

Nelder:2000:HHP

[Nel00b] J. A. Nelder. How helpful are packages in teaching statistics? Maths and Stats, 11(??):15–17, ???? 2000.

Nelder:2000:QLP

[Nel00c] J. A. Nelder. Quasi-likelihood and pseudo-likelihood are not the same thing. Journal of Applied Statistics, 27(8):1007–1011, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nelder:2000:TNO

[Nel00d] J. A. Nelder. There are no outliers in the stackloss data. Student, 3(??):211–216, ???? 2000.

Nelder:2000:ACT

[Nel00e] John A. Nelder. The analysis of contingency tables with one factor as the response: Round two. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):383–388, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/2681063.

Nelder:2001:C

[Nel01a] J. A. Nelder. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):209–211, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/2681095.

Nelder:2001:LES

[Nel01b] John A. Nelder. Letter to the Editor: Statistics in medical journals: some recent trends by Douglas G. Altman, Statistics in Medicine 2000; 19: 3275–3289. Statistics in Medicine, 20(14): 2205, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [Alt00].

Nelder:2004:F

[Nel04a] J. A. Nelder. Foreword. In Adams et al. [ACHS04c], pages xi– ?? ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Nelder:2004:SEH

[Nel04b] John A. Nelder. A statistical examination of the Hastings Rarities. In Adams et al. [ACHS04c], pages 215–234. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Nelder:2005:SQ

[Nel05] John Nelder. Synthesis and quality. Significance (Oxford, England), 2(2):75–77, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nelder:1984:MRP

[NF84] J. A. Nelder and E. L. Frome. Models for rates with Poisson errors. Biometrics, 40(4):1159–1162, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://

www.jstor.org/stable/2531169.

Nelder:1971:SSG

[NG71] J. A. Nelder and J. C. Gower. Statistical systems and generalpurpose languages. Bulletin of the International Statistical Institute, 44(??):296–301, ???? 1971.

Nelder:1994:DSQ

[NGL^{+}94] J. A. Nelder, T. Greenfield, H. J. Lenz, C. Chatfield, D. A. Preece, C. E. Lunneborg, M. C. Jones, J. Gower, R. A. Stone, M. C. Fessey, S. J. W. Evans, T. Lewis, A. S. C. Ehrenberg, D. J. Finney, A. M. Herzberg, P. Lovie, A. D. Lovie, R. J. Mackay, R. W. Oldford, I. W. Molenaar, P. C. Obrien, H. Rouanet, T. M. F. Smith, J. Tukey, M. Wise, and J. Zighera. Deconstructing statistical questions — discussion. Journal of the Royal Statistical Society Series A — Statistics in Society, 157(??):338–356, ???? 1994. ISSN 0035-9238.

Nelder:1994:RRR

[NGMS94] J. A. Nelder, E. I. George, U. E. Makov, and A. F. M. Smith. Reader reaction: a re-analysis of the pump-failure data [with response]. Scandinavian Journal of Statistics. Theory and Applications, 21(2):187–191, June 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https://www.jstor.

org/stable/4616310.

Nelder:1992:ESP

[NGN92] J. A. Nelder, Sander Greenland, and R. G. Newcombe. Estimating standardized parameters from generalized linear models. Statistics in Medicine, 11(4):559–560, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelder:1992:LE

[NKA92] J. A. Nelder, Stena Kettl, and Harold Ascher. Letters to the Editors: Pseudo-likelihood and quasi-likelihood. Applied Statistics, 41(3):595–600, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1991:GLMa

[NL91] J. A. Nelder and Y. Lee. Generalized linear models for the analysis of Taguchi-type experiments. Applied Stochastic Models and Data Analysis, 7(1):107–120, March 1991. ISSN 1099-0747 (print),

8755-0024 (electronic). URL https://onlinelibrary.wiley.
com/doi/10.1002/asm.3150070110.

Nelder:1992:LQL

[NL92] J. A. Nelder and Y. Lee. Likelihood, quasi-likelihood and pseudolikelihood: some comparisons. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):273–284, 1992. CO-DEN JSTBAJ. ISSN 0035-9246. URL http://links.jstor.

org/sici?sici=0035-9246(1992)54:1<273:LQAPSC>2.0.CO;2-
Y&origin=MSN.

Nelder:1994:DGL

[NL94] J. A. Nelder and Y. Lee. Double generalized linear models. Technical Report ??, Department of Mathematics, Imperial College, London, UK, ???? 1994.

Nelder:1995:CAF

[NL95] John A. Nelder and Peter W. Lane. The computer analysis of factorial experiments: In memorian — Frank Yates. The American Statistician, 49(4):382–385, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684580.

Nelder:1997:EQL

[NL97] J. A. Nelder and Y. Lee. Extended quasi-likelihood and estimating equations. In Selected Proceedings of the Symposium on Estimating Functions (Athens, GA, 1996), volume 32 of IMS Lecture Notes Monogr. Ser., pages 139–148. Institute of Mathematical Statistics, Hayward, CA, USA, 1997. ISBN 0-940600-44-7.

Nelder:2001:MAC

[NL01] J. A. Nelder and Y. Lee. Modelling and analysing correlated nonnormal data. Statistical Modelling, 1(??):3–16, ???? 2001.

Nelder:2004:DHG

[NL04a] J. A. Nelder and Y. Lee. Double hierarchical generalized linear models. Final paper listed in [ACHS04b], as submitted to Scandinavian Journal of Statistics, but apparently never published there, and not recorded in scandjstat.bib, ???? 2004.

Nelder:2004:RDG

[NL04b] J. A. Nelder and Y. Lee. Robust design via generalized linear models. J. Qual. Technol., 35(??):2–12, ???? 2004.

Nelder:1998:JMM

[NLB^{+}98] John A. Nelder, Youngjo Lee, Bo Bergman, Anders Hynén, A. Freek Huele, and Jan Engel. Joint modeling of mean and dispersion. Technometrics, 40(2):168–175, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270676.

Nelder:2008:LEW

[NLSS08] John A. Nelder, Viviana B. Lencina, Julio M. Singer, and Edward J. Stanek III. Letter to the Editors: What is the mixedmodels controversy? (with response). International Statistical Review = Revue Internationale de Statistique, 76(1):134–135, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919588.

Nijkamp:1985:MU

[NLW85] Peter Nijkamp, Helga Leitner, and Neil Wrigley, editors. Measuring the Unmeasurable, volume 22 of NATO ASI series. Series D. Behavioural and social sciences. Martinus Nijhoff, Dordrecht, The Netherlands, 1985. ISBN 90-247-3124-0. ix + 713 pp.

LCCN QA278.2 .N38 1983. URL http://www.gbv.de/dms/tib-
ub-hannover/01664381X.pdf.

Nelder:1956:SLD

[NM56a] J. A. Nelder and N. Moss. The spacing of lettuce in Dutch light frames and cold structures. Experimental Horticulture, (4):20–30, ???? 1956.

Nelder:1956:SLH

[NM56b] J. A. Nelder and N. Moss. The spacing of lettuce in heated glasshouses. Journal of Horticultural Science, 31(??):177–187, ???? 1956. CODEN JHSOFF. ISSN 2582-4899.

Nelder:1965:ESM

[NM65a] J. Nelder and R. Mead. Errata: “A simplex method for function minimization [MR3363409]. The Computer Journal, 8(1):27, 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Nelder:1965:SMF

[NM65b] J. A. Nelder and R. Mead. A simplex method for function minimization. The Computer Journal, 7(4):308–313, 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). See sixty-year retrospective [Gal24].

Nelder:1998:MIP

[NM98] J. A. Nelder and Z. Malik. Modified inverse polynomials as a response-surface model for magnetic flux density. Electronics Letters, 34(??):2252–2253, ???? 1998. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Nelder:1998:RSM

[NMD^{+}98] J. A. Nelder, Z. Malik, D. Dyck, R. Spence, and D. Lowther. Response-surface models using function values and gradient information, with application to the design of an electromagnetic device. In ????, editor, Engineering Design Conference ’98, pages 199–209. ????, ????, 1998.

Nelder:1998:IED

[NMS98] J. A. Nelder, Z. Malik, and H. Su. Informative experimental design for electronic circuits. Quality and Reliability Engineering International, 14(??):1–10, ???? 1998. CODEN QREIE5. ISSN 0748-8017 (print), 1099-1638 (electronic).

Nelder:1998:MPH

[NMT^{+}98] J. A. Nelder, Z. Malik, L. Tweedie, A. J. Smith, and R. Spence. Modelling for problem holders: the modellers guide. In R. Payne and P. Lane, editors, Compstat98 Proceedings in Computational Statistics, Bristol, pages 185–186. ????, ????, 1998.

Nelder:1975:DSS

[NP75] J. A. Nelder and R. W. Payne. Data structures in statistical computing. In Anonymous [Ano75], pages 199–207. ISBN ???? LCCN QH323.5 .I55 1976.

Nelder:1987:EQL

[NP87] J. A. Nelder and D. Pregibon. An extended quasi-likelihood function. Biometrika, 74(2):221–232, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336136.

Nelder:1991:GCE

[NP91] John A. Nelder and Roger W. Payne. GENSTAT as a computing environment. In Buja and Tukey [BT91], pages 133–138. ISBN 0-387-97633-7. LCCN QA276.4 .C59 1991.

Nelder:1974:STR

[NPT^{+}74] J. A. Nelder, G. P. Patil, J. L. Teugels, T. Dalenius, Saw Swee-Hock, and J. S. Siegel. Statistical training and research/formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 42(1):79–87, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402687.

Nelder:1997:LE

[NR97] J. A. Nelder and G. J. S. Ross. Letters to the Editors: Functional marginality is important. Applied Statistics, 46(2):281– 286, 1997. CODEN APSTAG. ISSN 0035-9254 (print), 1467-

9876 (electronic). URL http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0035-9254&src=ard&aid=066&iid=
2&vid=46.

Nelder:1991:GLMb

[NRCC91] J. A. Nelder, D. Ruppert, N. Cressie, and R. J. Carroll. Generalized linear models for enzyme-kinetic data. Biometrics, 47(4): 1605–1615, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532412.

Nelder:1974:LE

[NS74] J. A. Nelder and David A. Swick. Letters to the editors. Applied Statistics, 23(2):232–233, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nazareth:2002:GLV

[NT02] Larry Nazareth and Paul Tseng. Gilding the lily: a variant of the Nelder–Mead algorithm based on golden-section search. Computational optimization and applications, 22(1):133–144, 2002. CO-DEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Nelder:1997:CTG

[NV97] J. A. Nelder and R. J. Verrall. Credibility theory and generalized linear models. Astin Bulletin, 27(??):71–82, ???? 1997.

Nelder:1972:GLM

[NW72] J. A. Nelder and R. W. M. Wedderburn. Generalized linear models. Journal of the Royal Statistical Society. Series A (General), 135(3):370–384, ???? 1972. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/2344614.

Nelder:1986:FEG

[NW86] J. A. Nelder and D. E. Wolstenholme. A front-end for GLIM. In Boardman and Stefanski [BS86], pages 113–117. LCCN QA276.4 .S95 1986.

Nawaz:2015:JNH

[NZE^{+}15] M. Nawaz, A. Zeeshan, R. Ellahi, S. Abbasbandy, and Saman Rashidi. Joules and Newtonian heating effects on stagnation point flow over a stretching surface by means of genetic algorithm and Nelder–Mead method. International Journal of Numerical Methods for Heat & Fluid Flow, 25(3):665–684, 2015.

Olsson:1975:NMS

[ON75] Donald M. Olsson and Lloyd S. Nelson. The Nelder–Mead simplex procedure for function minimization. Technometrics, 17(1):45– 51, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1267998.

Ouria:2009:ANM

[OT09] Ahad Ouria and Mohammad M. Toufigh. Application of Nelder– Mead simplex method for unconfined seepage problems. Applied Mathematical Modelling, 33(9):3589–3598, 2009. CODEN AM-MODL. ISSN 0307-904X (print), 1872-8480 (electronic).

Pawitan:2004:LPC

[Paw04] Y. Pawitan. Likelihood perspectives in the consensus and controversies of statistical modelling and inference. In Adams et al. [ACHS04c], pages 23–52. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Payne:2004:ADS

[Pay04] R. Payne. Algorithms, data structures and languages: the computational ingredients for innovative analysis. In Adams et al. [ACHS04c], pages 95–118. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Puhan:2026:OPV

[PB26] Janez Puhan and Árpád Bűrmen. On the optimal parameter values of the Nelder–Mead simplex algorithm. Journal of Computational and Applied Mathematics, 484:Paper No. 117521, 18, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Price:2002:CVN

[PCB02] C. J. Price, I. D. Coope, and D. Byatt. A convergent variant of the Nelder–Mead algorithm. Journal of Optimization Theory and Applications, 113(1):5–19, 2002. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic).

Peren:2024:NON

[Per24] Franz W. Peren. Nonlinear optimization: the Nelder–Mead simplex search procedure. In Operations research and management— quantitative methods for planning and decision-making in business and economics, Springer Texts Bus. Econ., pages 187–199. Springer, Cham, 2024. ISBN 3-031-47205-5; 3-031-47206-3.

Peren:2026:NON

[Per26] Franz W. Peren. Nonlinear optimization: the Nelder–Mead simplex search procedure. In Operations research and management— quantitative methods for planning and decision-making in business and economics, Springer Texts Bus. Econ., pages 187–199. Springer, Cham, 2026. ISBN 3-032-08077-0; 3-032-08078-9.

Phelps:1987:IAI

[Phe87] Bob Phelps, editor. Interactions in Artificial Intelligence and Statistical Methods. Unicom applied information technology reports series. Gower Technical Press, Aldershot, UK, 1987. ISBN 0-291-39743-3 (paperback). xvii + 187 pp. LCCN Q334 .I548 1987.

Plackett:1984:BRB

[Pla84] R. L. Plackett. Book review: Generalized Linear Models, by P. McCullagh; J. A. Nelder. Journal of the Royal Statistical Society. Series A (General), 147(3):526–527, ???? 1984. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.jstor.org/stable/2981609.

Pandey:2023:PNI

[PP23] Vivek Pandey and Sudhir K. Pandey. PY-Nodes: an ab-initio Python code for searching nodes in a material using Nelder–Mead’s simplex approach. Computer Physics Communications, 283(??): ??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002892.

Pregibon:1984:RPM

[Pre84] Daryl Pregibon. Review: P. McCullagh, J. A. Nelder, Generalized Linear Models. Annals of Statistics, 12(4):1589–1596, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346819.

Price:2021:MNM

[Pri21] C. J. Price. A modified Nelder–Mead barrier method for constrained optimization. Numerical Algebra, Control and Optimization, 11(4):613–631, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http://aimsciences.org/

article/doi/10.3934/naco.2020058.

Park:1999:SPM

[PV99] Sunh H. Park and G. Geoffrey Vining, editors. Statistical Process Monitoring and Optimization, volume 160. Marcel Dekker, Inc., New York, NY, USA, 1999. ISBN 0-8247-6007-7 (hardcover). xii + 489 pp. LCCN TS156.8 .S7537 2000.

Ripley:2004:SAL

[Rip04] B. D. Ripley. Selecting amongst large classes of models. In Adams et al. [ACHS04c], pages 155–170. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Robinson:2007:BRBb

[Rob07] Andrew Robinson. Book review: Generalized Linear Models with Random Effects: Unified Analysis via H-Likelihood, by Y. Lee; J. A. Nelder; Y. Pawitan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):855–856, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/4623207.

Rodriguez:1995:CJN

[RTW95] Robert Rodriguez, Randall Tobias, and Russell Wolfinger. Comments on J. A. Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):97–101, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF00143938. See [Nel94d].

Ren:2013:HED

[RWJ13] Aihong Ren, Yuping Wang, and Fei Jia. A hybrid estimation of distribution algorithm and Nelder–Mead simplex method for solving a class of nonlinear bilevel programming problems. Journal of Applied Mathematics, pages Art. ID 378568, 10, 2013. ISSN 1110-757X (print), 1687-0042 (electronic).

Schuenemeyer:1992:BRB

[Sch92] John H. Schuenemeyer. Book review: Generalized Linear Models by P. McCullagh; J. A. Nelder. Technometrics, 34(2):224, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269238.

Searle:1995:CJN

[Sea95] Shayle R. Searle. Comments on J. A. Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):103– 107, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-

1375 (electronic). URL http://link.springer.com/article/
10.1007/BF00143939. See [Nel94d].

Suhail:2024:IDH

[SEABZ24] Enas Suhail, Mahmoud El-Alem, Omar Bazighifan, and Ahmed Zekri. Improving diversification by a hybrid bat-Nelder–Mead algorithm and DDE for rapid convergence to solve global optimization. AIMS Mathematics, 9(12):35655–35677, 2024. ISSN 2473-6988.

Selvan:2015:NIC

[Sel15] Suviseshamuthu Easter Selvan. Nonsmooth ICA contrast minimization using a Riemannian Nelder–Mead method. IEEE Transactions on Neural Networks and Learning Systems, 26(1):177–183, 2015.

Senn:2003:CJN

[Sen03] Stephen Senn. A conversation with John Nelder. Statistical Science, 18(1):118–131, February 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1056397489.

Senn:2004:JNG

[Sen04] S. Senn. John Nelder: From general balance to generalised models (both linear and hierarchical). In Adams et al. [ACHS04c], pages 1–12. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Senn:2019:JAN

[Sen19] Stephen Senn. John Ashworth Nelder. 8 October 1924–7 August 2010. Biographical Memoirs of Fellows of the Royal Society, 67:307–326, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0013.

Sarakhsi:2016:NHA

[SGK16] M. Khojaste Sarakhsi, S. M. T. Fatemi Ghomi, and B. Karimi. A new hybrid algorithm of scatter search and Nelder–Mead algorithms to optimize joint economic lot sizing problem. Journal of Computational and Applied Mathematics, 292(??):387–401, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042715003891.

Smith:1999:VTM

[SMNS99] A. J. Smith, Z. Malik, J. A. Nelder, and R. Spence. Visualisation tools for model making. In Herwig Friedl, Andrea Berghold, and Göran Kauermann, editors, Proceedings of the 14th International Workshop on Statistical Modelling. Graz, Austria, July 19–23, 1999, page ?? Statistical Modelling Soci-

ety, ????, 1999. URL https://www.statmod.org/workshops_
archive_proceedings_1999.htm.

Smith:2001:VIM

[SMNS01] Andrew J. Smith, Zahid Malik, John A. Nelder, and Robert Spence. A visual interface for model-fitting. Quality and Reliability Engineering International, 17(2):85–91, March/April 2001. CODEN QREIE5. ISSN 0748-8017 (print), 1099-1638

(electronic). URL https://onlinelibrary.wiley.com/doi/10.
1002/qre.384.

Smith:2001:SSM

[SNB^{+}01] Andrew J. Smith, John Nelder, Andreas Buja, Zahid Malik, Lisa Tweedie, and Robert Spence. Sampling schemes for model visualization. Journal of Computational and Graphical Statistics, 10(3): 545–554, 2001. ISSN 1061-8600 (print), 1537-2715 (electronic).

Singer:2001:CAN

[SS01] Sanja Singer and Saša Singer. Complexity analysis of Nelder– Mead search iterations. In Applied mathematics and computation (Dubrovnik, 1999), pages 185–196. Univ. Zagreb Dep. Math., Zagreb, 2001. ISBN 953-6076-70-5.

Singer:2004:EIN

[SS04] Saša Singer and Sanja Singer. Efficient implementation of the Nelder–Mead search algorithm. ANACM. Applied Numerical Analysis and Computational Mathematics, 1(3):524–534, 2004. ISSN 1611-8170 (print), 1611-8189 (electronic).

Streitberg:1988:NES

[Str88] Bernd Streitberg. On the nonexistence of expert systems: Critical remarks on artificial intelligence in statistics. Statistical Software Newsletter, 14(??):55–62, ???? 1988. ISSN 0173-5896.

Turksen:2013:EES

[TA13] Özlem Türkşen and Ayşen Apaydin. Estimating the earthquake source parameters: simulated annealing versus Nelder–Mead simplex algorithm. Communications de la Faculté des Sciences de l’Université d’Ankara. Séries A1. Mathematics and Statistics. Communications. Faculty of Sciences. University of Ankara, 62 (2):53–66, 2013. ISSN 1303-5991.

Tekile:2021:CEM

[TFB21] Hailemariam Abebe Tekile, Michele Fedrizzi, and Matteo Brunelli. Constrained eigenvalue minimization of incomplete pairwise comparison matrices by Nelder–Mead algorithm. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/222.

Tomick:1995:CNM

[Tom95] John James Tomick. On convergence of the Nelder–Mead simplex algorithm for unconstrained stochastic optimization. Thesis (Ph.D.), The Pennsylvania State University, State College,

PA, USA, 1995. 188 pp. URL https://www.proquest.com/
pqdtglobal/docview/304210989.

Takenaga:2023:PIN

[TOO23] Shintaro Takenaga, Yoshihiko Ozaki, and Masaki Onishi. Practical initialization of the Nelder–Mead method for computationally expensive optimization problems. Optimization Letters, 17(2): 283–297, 2023. ISSN 1862-4472 (print), 1862-4480 (electronic).

Ulin:1958:CNM

[Uli58] Bengt Ulin. On a conjecture of Nelder in mathematical statistics. Compositio Mathematica, 13:148–149, 1958. CODEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic).

VanEeuwijk:1995:TDS

[Van95] F. A. Van Eeuwijk. On the tenability of distinctions: some comments on ‘The statistics of linear models: back to basics’ by John Nelder. Statistics and Computing, 5(2):93–95, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF00143937. See [Nel94d].

Wakefield:2004:NLR

[Wak04] J. C. Wakefield. Non-linear regression modelling and inference. In Adams et al. [ACHS04c], pages 119–154. ISBN 1-281-86647-4, 1-86094-463-9 (hardcover), 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Watson:1986:BRG

[Wat86] G. S. Watson. Book review: Generalized Linear Models (P. Mccullagh and J. A. Nelder). SIAM Review, 28(1):128–130, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wessing:2019:PIC

[Wes19] Simon Wessing. Proper initialization is crucial for the Nelder– Mead simplex search. Optimization Letters, 13(4):847–856, 2019. ISSN 1862-4472 (print), 1862-4480 (electronic).

Williams:1984:BRG

[Wil84] D. A. Williams. Book review: Generalized Linear Models, by P. McCullagh, J. A. Nelder. Biometrics, 40(2):566, June 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531415.

Wirjanto:1997:CFD

[Wir97] T. S. Wirjanto. Comment on A. F. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):109– 113, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900136. See [Des97a].

Wolstenholme:1986:FEG

[WN86] D. E. Wolstenholme and J. A. Nelder. A front end for GLIM. In Expert Systems in Statistics: Selected Papers from a Workshop, organized by the Working Group “Computational Statistics” of the German Region of the International Biometric Society, December 6–7, 1985, Aachen, F.R.G. [Hau86], pages 155–177. ISBN 0-89574-235-7 (paperback), 3-437-40171-8 (paperback). LCCN QA276.4 .E925 1986.

Wolstenholme:1988:GKB

[WON88] David E. Wolstenholme, Carol M. O’Brien, and John A. Nelder. GLIMPSE: a knowledge-based front end for statistical analysis. Knowledge-Based Systems, 1(3):173–178, June 1988. CODEN KN-SYET. ISSN 0950-7051 (print), 1872-7409 (electronic).

Wright:2012:NMO

[Wri12] Margaret H. Wright. Nelder, Mead, and the other simplex method. Documenta mathematica, pages 271–276, 2012. ISSN 1431-0643 (print), 1431-0643 (electronic).

Wyers:2013:BDN

[WSKF13] Eric J. Wyers, Michael B. Steer, C. T. Kelley, and Paul D. Franzon. A bounded and discretized Nelder–Mead algorithm suitable for RFIC calibration. IEEE Transactions on Circuits and Systems I: Regular Papers, 60(7):1787–1799, 2013. ISSN 1549-8328 (print), 1558-0806 (electronic).

Wang:2015:FBS

[WYWW15] Ze Wang, Limin Yang, Chi Man Wong, and Feng Wan. Fast basis searching method of adaptive Fourier decomposition based on Nelder–Mead algorithm for ECG signals. In Advances in neural networks—ISNN 2015, volume 9377 of Lecture Notes in Comput. Sci., pages 305–314. Springer, Cham, 2015. ISBN 3-319-25393-X; 3-319-25392-1.

Xiao:2014:MDB

[XD14] Hong Feng Xiao and Ji An Duan. Multi-direction-based Nelder– Mead method. Optimization, 63(7):1005–1026, 2014. CODEN OPTZDQ. ISSN 0233-1934, 0323-3898.

Yang:2024:JIJ

[YCH24] Yuxin Yang, Xiaoming Chen, and Yinhe Han. JANM-IK: Jacobian argumented Nelder–Mead algorithm for inverse kinematics and its hardware acceleration. IEEE Computer Architecture Letters, 23 (1):45–48, January/June 2024. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Young:1976:LE

[YYB^{+}76] Kan Hua Young, Lih Ying Young, A. E. Brandt, C. E. Werts, D. A. Rock, R. L. Linn, K. C. Joreskog, R. W. Farebrother, Oliver D. Anderson, Federico J. O’Reilly, J. A. Nelder, Nathan Mantel, Mark J. Nicolich, and John Neter. Letters to the Editor. The American Statistician, 30(2):101–104, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683805.

Zhang:2007:SMU

[ZHZ^{+}07] Kai Zhang, Ming You Huang, Ran Zhang, Gang Wang, and Jin Yue Gao. Simulation measurement of ultrafine nanoparticle size: Nelder–Mead simplex method. Gongcheng Shuxue Xuebao, 24(6):977–986, 2007. ISSN 1005-3085.

Zhu:2022:IHH

[ZZP^{+}22] Cheng Zhu, Yong Zhang, Xuhua Pan, Qi Chen, and Qingyu Fu. Improved Harris hawks optimization algorithm based on quantum correction and Nelder–Mead simplex method. Mathematical Biosciences and Engineering. MBE, 19(8):7606–7648, 2022.